### v2

# Emmanuel Letellier Université Paris Cité, IMJ-PRG CNRS UMR 7586

Abstract
In this paper we study the mixed Poincaré polynomials of generic parabolic $\operatorname{PGL}_{2}(\mathbb{C})$-character varieties with coefficients in some local systems arising from the conjugacy class of $\operatorname{PGL}_{2}(\mathbb{C})$ which has a non-connected stabiliser. We give a conjectural formula that we prove to be true under the Euler specialisation. We then prove that these conjectured formulas interpolate the structure coefficients of the two following based rings:
$\left(\mathcal{C}(\operatorname{PGL}_{2}(\mathbb{F}_{q})),Loc(\operatorname{PGL}_{2}),*\right),\hskip 28.45274pt\left(\mathcal{C}(\operatorname{SL}_{2}(\mathbb{F}_{q})),CS(\operatorname{SL}_{2}),\cdot\right)$
where for a group $H$, $\mathcal{C}(H)$ denotes the space of complex valued class functions on $H$, $Loc(\operatorname{PGL}_{2})$ denotes the basis of characteristic functions of intermediate extensions of equivariant local systems on conjugacy classes of $\operatorname{PGL}_{2}$ and $CS(\operatorname{SL}_{2})$ the basis of characteristic functions of Lusztig’s character-sheaves on $\operatorname{SL}_{2}$. Our result reminds us of a non-abelian Fourier transform.
1 Introduction
Given a finite group $H$ we denote by $\mathcal{C}(H)$ the $\mathbb{C}$-vector space of $\mathbb{C}$-valued class functions on $H$. It is equipped with the basis $Cl(H)$ of characteristic functions of the conjugacy classes of $H$ and the basis $\widehat{H}$ of irreducible characters of $H$. We then consider the two rings
[…]
where * is the convolution product and $\cdot$ the pointwise multiplication.
The first ring can be identified with the center of the group algebra of $H$ and the second one is the character ring of $H$.
These rings are determined by their structure coefficients
$\left\langle 1_{C_{1}}*1_{C_{2}},1_{C_{3}}\right\rangle_{H},\hskip 28.45274pt\left\langle\chi_{1}\otimes\chi_{2},\chi_{3}\right\rangle_{H}$
where $\langle\,,\,\rangle_{H}$ is the usual inner product on $\mathcal{C}(H)$ making the basis $\widehat{H}$ orthonormal and where $C_{1},C_{2},C_{3}$ run over the set of conjugacy classes of $H$ and $\chi_{1},\chi_{2},\chi_{3}$ run over $\widehat{H}$.
When $H$ is abelian, these two rings are isomorphic, namely there is a (non-canonical) automorphism of $\mathcal{C}(H)$ mapping $Cl(H)$ to $\widehat{H}$ and transforming * into $\cdot$ (which can be thought as a Fourier transform).
We are interested in the case where $H$ is the group $G(\mathbb{F}_{q})$ of rational points of a connected reductive group $G$ defined over a finite field $\mathbb{F}_{q}$.
For that purpose we need base other than $Cl(G^{F})$ and $\widehat{G^{F}}$ :
Denote by $F:G\rightarrow G$ the geometric Frobenius (then $G^{F}=G(\mathbb{F}_{q})$).
Instead of the basis $Cl(G^{F})$ we consider the basis $Loc(G)^{F}$ formed by the characteristic functions of the intermediate extensions of the irreducible $F$-equivariant $G$-equivariant local systems on the $F$-stable conjugacy classes of $G$, and instead of the basis $\widehat{G^{F}}$ we consider the basis $CS(G)^{F}$ formed by the characteristic functions of the $F$-equivariant Lusztig’s character-sheaves on $G$ (which are closely related to irreducible characters and coincide often with them).
We are interested in the relationship between the structure coefficients of the two based rings
$\mathcal{A}_{G}=\left(\mathcal{C}(G^{F}),Loc(G)^{F},*\right),\hskip 28.45274pt\mathcal{B}_{G^{*}}=\left(\mathcal{C}(G^{*}{{}^{F}}),CS(G^{*})^{F},\cdot\right)$
where $(G^{*},F)$ is the dual group of $(G,F)$ in the sense of Deligne-Lusztig [8].
1.1 Review of the case $G=\mathrm{GL}_{n}$
In this case $G=G^{*}$ and the basis $CS(G)^{F}$ coincides with $\widehat{G^{F}}$. The stabilisers of the elements of $G$ being all connected, the irreducible $G$-equivariant local systems on the conjugacy classes are the constant sheaf. Therefore, the elements of $Loc(G)^{F}$ and $Cl(G^{F})$ attached to semisimple conjugacy classes coincide.
Given a generic $k$-tuple $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ of conjugacy classes of $G(\mathbb{C})$ we consider the quotient stack (possibly empty)
$\mathcal{M}_{\overline{\mathcal{C}}}=\left[\{(x_{1},\dots,x_{k})\in\overline{\mathcal{C}_{1}}\times\cdots\times\overline{\mathcal{C}_{k}}\,|\,x_{1}\cdots x_{k}=1\}/G\right].$
The genericity assumption ensures that the diagonal action of $G$ induces a free action of $\operatorname{PGL}_{n}$. We could have decided to take the quotient by $\operatorname{PGL}_{n}$ instead of $G=\operatorname{GL}_{n}$ but this would require to correct some formulas (like Formula (1.1.2) by a factor $q-1$.
The stack $\mathcal{M}_{\overline{\mathcal{C}}}$, called a $G$-character stack, is the stack of $G$-local systems on the Riemann sphere minus $k$ points $a_{1},\dots,a_{k}$ with local monodromies around $a_{i}$ in $\overline{\mathcal{C}_{i}}$.
We consider the compactly supported intersection cohomology $IH_{c}^{i}(\mathcal{M}_{\overline{\mathcal{C}}},\mathbb{C})$ of $\mathcal{M}_{\overline{\mathcal{C}}}$.
It is equipped with a weight filtration (increasing) $W^{i}_{\bullet}$ from which we define the mixed Poincaré series
$IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)=\sum_{i}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}.$
Our conjugacy classes are defined over a subring $R$ of $\mathbb{C}$ which is finitely generated as a $\mathbb{Z}$-algebra. We can choose this subring $R$ such that for any finite field $\mathbb{F}_{q}$ and any ring homomorphism $\varphi:R\rightarrow\mathbb{F}_{q}$, the conjugacy classes of $G(\overline{\mathbb{F}}_{q})$ obtained from $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ by base change are of same Jordan type and form a generic $k$-tuple. To ease the notation we denote them again by $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$.
Then we have [18, Theorem 4.13]
$E(\mathcal{M}_{\overline{\mathcal{C}}};q):$ $=IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,-1)$ (1.1.1)
$=\left\langle{\bf Y}_{\mathcal{C}_{1}}*\cdots*{\bf Y}_{\mathcal{C}_{k}},1_{\{1\}}\right\rangle_{G(\mathbb{F}_{q})}$ (1.1.2)
where ${\bf Y}_{\mathcal{C}}:G(\mathbb{F}_{q})\rightarrow\overline{{\mathbb{Q}}}_{\ell}\simeq\mathbb{C}$ denotes the characteristic function of the intersection cohomology complex of $\overline{\mathcal{C}}$ (in the $\ell$-adic setting).
In [18, Conjecture 4.5] we give a conjecture formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ in terms of Macdonald symmetric functions. When the conjugacy classes are semisimple, this was first conjectured in [11] (see §2.2 for a review in the semisimple case). The conjecture is known to be true in some non-trivial examples where we can compute explicitly the weight filtration (see [11, §1.5.3]).
This conjectural formula is proved after the specialisation $t\mapsto-1$ (see [18, Theorem 4.8] and [11, Theorem 1.2.3] in the semisimple case). It is also proved after the specialisation $q\mapsto 1$ which gives the Poincaré series. In the semisimple case this is due to by A. Mellit [27] who followed a strategy used by O. Schiffmann to compute the Poincaré polynomial of the moduli space of semistable Higgs bundles over a smooth projective curve [30]. In the general case, this is due to M. Ballandras [2].
We define the "pure part" of $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ as
$PIH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q)=\sum_{i}{\rm dim}\,\left(W^{i}_{i}/W^{i}_{i-1}\right)q^{i/2}.$
The conjectural formula for $IH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q,t)$ implies a conjectural formula for $PIH_{c}(\mathcal{M}_{\overline{\mathcal{C}}};q)$ in terms of Hall-Littlewood symmetric functions. In [19, Theorem 6.10.1] we prove that the conjectured formula for the pure part coincides with the inner product
$\left\langle\chi_{1}\otimes\cdots\otimes\chi_{k},1\right\rangle_{G(\mathbb{F}_{q})}$
where $(\chi_{1},\dots,\chi_{k})$ is a generic $k$-tuple of irreducible characters of $G(\mathbb{F}_{q})$ of same Jordan type (in the sense of Lusztig) as the conjugacy classes $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$.
Therefore, it follows from the main conjecture for the mixed Poincaré series of character stacks that the latter interpolate (generically) the structure coefficients of the two rings $\mathcal{A}_{\operatorname{GL}_{n}}$ and $\mathcal{B}_{\operatorname{GL}_{n}}$.
The non-generic situation is more complicated and will not be discussed in this paper. However we refer to [28][29] for results in the non-generic semisimple case.
Finally to conclude this part on the $G=\operatorname{GL}_{n}$ case, notice that there is a notion of character-sheaves on the Lie algebra of $G$ due to Lusztig [24] (see also [17]). The Lie algebra analogues $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ of these two rings $\mathcal{A}$ and $\mathcal{B}$ are then isomorphic by an arithmetic Fourier transform (the character-sheaves are precisely the Deligne-Fourier transforms of the intersection cohomology complexes on the Zariski closures of the adjoint orbits of the Lie algebra). The generic structure coefficients of $\mathcal{A}^{Lie}$ and $\mathcal{B}^{Lie}$ are thus equal. Moreover we can prove that the generic structure coefficients of $\mathcal{B}^{Lie}$ and $\mathcal{B}$ are also equal [19, Theorem 6.9.1] and coincide [11][19] with the Poincaré series (for intersection cohomology) of
$\mathcal{Q}=\left[\{(x_{1},\dots,x_{k})\in\overline{\mathcal{O}_{1}}\times\dots\times\overline{\mathcal{O}_{k}}\,|\,x_{1}+\cdots+x_{k}=0\}/G\right]$
where $(\mathcal{O}_{1},\dots,\mathcal{O}_{k})$ runs over the set of generic $k$-tuple of adjoint orbits of ${\rm gl}_{n}(\mathbb{C})$ (unlike the multiplicative case, generic $k$-tuples of adjoint orbits of given Jordan types do not always exist).
Finally, unlike character stacks, the intersection cohomology of $\mathcal{Q}$ is pure [11, Proposition 2.2.6][19, Theorem 7.3.2] and its Poincaré series is conjectured to be the pure part of the mixed Poincaré series of character stacks. This latter conjecture is also a consequence of the conjectural formula for the mixed Poincaré series of generic character stacks.
1.2 The case where $G=\mathrm{PGL}_{2}$
The aim of this paper is to verify for a group $G\neq G^{*}$ the assumption that the mixed Poincaré series of $G$-character-stack interpolate structure coefficients of the two rings $\mathcal{A}_{G}$ and $\mathcal{B}_{G^{*}}$. We are especially interested in the situation where local monodromies are in conjugacy classes which carry non-trivial $G$-equivariant local systems.
We now assume that $G=\operatorname{PGL}_{2}$.
When $K$ is an algebraically closed field of characteristic $\neq 2$, there is only one conjugacy class of $G(K)$ which has a disconnected centraliser. This is the $G(K)$-conjugacy class $\mathcal{C}_{-1}$ of the image in $\operatorname{PGL}_{2}(K)$ of the diagonal matrix with coefficients $1$ and $-1$. There are thus two irreducible (up to isomorphism) $G(K)$-equivariant local systems on $\mathcal{C}_{-1}$ which are naturally parametrised by the set $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{1,\epsilon\}$ of irreducible characters of the group of connected components of the centraliser. We will denote by $\mathcal{L}_{1}$ and $\mathcal{L}_{\epsilon}$ the trivial and non-trivial local systems on $\mathcal{C}_{-1}$ (when the characteristic of $K$ is positive, the local systems considered are in the $\ell$-adic setting).
We will call $\mathcal{C}_{-1}$ the degenerate conjugacy class. The other conjugacy classes (non-degenerate) have all a connected centraliser and so the constant sheaf is the only irreducible $G(K)$-equivariant local system on them.
$\operatorname{PGL}_{2}$-character varieties
Consider a generic $k$-tuple $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ of semisimple regular conjugacy classes of $G(\mathbb{C})$ (see Definition 3.2.3) and consider the $G(\mathbb{C})$-character stack
$\mathcal{M}_{\mathcal{C}}=\mathcal{M}_{\mathcal{C}}(\mathbb{C}):=[\{(x_{1},\dots,x_{k})\in\mathcal{C}_{1}\times\cdots\times\mathcal{C}_{k}\,|\,x_{1}\cdots x_{k}=1\}/G].$
Because of the genericity condition this stack is in fact a smooth algebraic variety of dimension $62k-$ 6 (see Theorem 3.2.6).
Choose on each $\mathcal{C}_{i}$ an irreducible $G$-equivariant local system $\mathcal{L}_{i}$ (it is the trivial local system if $\mathcal{C}_{i}$ is non-degenerate and can be either $\mathcal{L}_{1}$ or $\mathcal{L}_{\epsilon}$ if $\mathcal{C}_{i}=\mathcal{C}_{-1}$). This defines a local system $\mathcal{L}_{\mathcal{C}}$ on $\mathcal{M}_{\mathcal{C}}$.
We consider the mixed Poincaré polynomial
$H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}};q,t)=\sum_{i}{\rm dim}\left(W^{i}_{r}/W^{i}_{r-1}\right)q^{r/2}t^{i}$
where $W^{i}_{\bullet}$ is the weight filtration on $H_{c}^{i}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}})$.
In this paper we give an explicit conjectural formula for $H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}};q,t)$ (see Conjecture 3.4.3 for the non-degenerate case and Conjecture 3.4.6 when at least one conjugacy class is degenerate) and we prove that it is true after the specialisation $t\mapsto-1$ (see Theorem 3.4.2 and Theorem 3.4.5).
Our conjectural formula for $H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{L}_{\mathcal{C}};q,t)$ is a consequence of the conjectural formula for the mixed Poincaré polynomials of $\operatorname{GL}_{n}$-character varieties for which we have many evidences.
Langlands duality over finite fields (Lusztig) when $G=\operatorname{PGL}_{2}$
Assume that $K=\overline{\mathbb{F}}_{q}$ with odd $q$ and that $F:G\rightarrow G$ is the standard Frobenius ($G(\mathbb{F}_{q})=G^{F}$).
We now explain the correspondence between $F$-stable local systems on $F$-stable semisimple conjugacy classes of $\operatorname{PGL}_{2}(K)$ and $F$-stable (semisimple) character-sheaves on $\operatorname{SL}_{2}(K)$.
For an $F$-stable $G$-equivariant local system $\mathcal{E}$ on an $F$-stable semisimple conjugacy class $\mathcal{C}$ of $G$, we denote by ${\bf Y}_{(\mathcal{C},\mathcal{E})}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$ the characteristic function of $\mathcal{L}_{\epsilon}$ defined by
${\bf Y}_{(\mathcal{C},\mathcal{E})}(x)=\operatorname{Tr}(Frob,\mathcal{E}_{x}).$
(1) If $\mathcal{E}=\overline{{\mathbb{Q}}}_{\ell}$, then ${\bf Y}_{(\mathcal{C},\overline{{\mathbb{Q}}}_{\ell})}$ is the function $1_{\mathcal{C}^{F}}$ that takes the value $1$ on $\mathcal{C}^{F}$ and 00 elsewhere.
(2) If $(\mathcal{C},\mathcal{E})=(\mathcal{C}_{-1},\mathcal{L})$, then
${\bf Y}_{(\mathcal{C},\mathcal{E})}=1_{\mathcal{O}_{e}}-1_{\mathcal{O}_{\sigma}}$
where
$(\mathcal{C}_{-1})^{F}=\mathcal{O}_{e}\sqcup\mathcal{O}_{\sigma}$
is the decomposition into $G^{F}$-conjugacy classes (see Remark 5.1.2).
Denote by $T$ and $T^{*}$ the maximal tori respectively of $\operatorname{PGL}_{2}$ and $\operatorname{SL}_{2}$ of diagonal matrices.
We have a non-canonical bijection
$T^{F}\simeq\mathbb{F}_{q}^{*}\overset{\sim}{\longrightarrow}\widehat{\mathbb{F}_{q}^{*}}\simeq\widehat{T^{*}{{}^{F}}}$
under which $-1$, the non-trivial square of $1$, corresponds to the non-trivial square $\alpha_{o}$ of the identity character of $T^{*}{{}^{F}}$. The character $\alpha_{o}$ takes the value $1$ at squares elements and $-1$ at non-square elements.
Now the linear characters of $T^{*}{{}^{F}}$ are in bijection with the isomorphism classes of the $F$-stable Kummer local systems on $T^{*}$ (see Proposition 4.3.1).
We thus have a bijection $x\mapsto\mathcal{E}_{x}$ from $\mathbb{F}_{q}^{*}\simeq T^{F}$ to the set of isomorphism classes of $F$-stable Kummer local systems on $T^{*}$.
Denote by $\mathcal{I}_{T^{*}}^{G^{*}}:\mathcal{M}(T^{*};F)\rightarrow\mathcal{M}(G^{*};F)$ the geometric induction functor between categories of $F$-equivariant perverse sheaves (see §4.4).
Let $\mathcal{E}_{x}$ be an $F$-stable Kummer local system on $T^{*}$. We have the following two situations :
(1) If $x\neq-1,1$, then $\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})$ is an $F$-stable irreducible $G^{*}$-equivariant perverse sheaf on $G^{*}$ which is an example of a character-sheaf (its characteristic function is an irreducible character of $G^{*}{{}^{F}}$).
(2) If $x=-1$, then $\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})$ is an $F$-stable $G^{*}$-equivariant semisimple perverse sheaf whose simple direct factors are naturally paramerised by $\widehat{\mathbb{Z}/2\mathbb{Z}}=\{1,\epsilon\}$ :
$\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})=\mathcal{X}_{1}[3]\oplus\mathcal{X}_{\epsilon}[3].$
The perverse sheaves $\mathcal{X}_{1}[3]$ and $\mathcal{X}_{\epsilon}[3]$ are other examples of character-sheaves. However unlike the ones above their characteristic functions are not irreducible characters (see the end of §4.5).
We thus have a bijection between irreducible equivariant local systems on semisimple conjugcy classes of $G$ and certain character sheaves on $G^{*}$ :
$\Phi:\left\{(\mathcal{C}_{x},\overline{{\mathbb{Q}}}_{\ell})\,|\,x\neq 1,-1\right\}\sqcup\{(\mathcal{C}_{-1},\mathcal{L}_{r})\,|\,r=1,\epsilon\}\simeq\left\{\mathcal{I}_{T^{*}}^{G^{*}}(\mathcal{E}_{x})\,|\,x\neq 1,-1\right\}\sqcup\left\{\mathcal{X}_{r}[3]\,|\,r=1,\epsilon\right\}$
where $\mathcal{C}_{x}$ is the (semisimple) $G$-conjugacy class of $x\in\mathbb{F}_{q}^{*}\simeq T^{F}$.
Notice that on the left hand side we are missing the unipotent conjugacy classes and also the non-split ones (i.e. the ones with eigenvalues in $\mathbb{F}_{q^{2}}\backslash\mathbb{F}_{q}$) and on the right hand side we are missing the unipotent character-sheaves (the simple direct factors of $\mathcal{I}_{T^{*}}^{G^{*}}(\overline{{\mathbb{Q}}}_{\ell})$ whose characteristic functions are the irreducible unipotent characters of $G^{*}{{}^{F}}$) and the two cuspidal character-sheaves (see §4.5).
Main result
Let $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ be a generic $k$-tuple of semisimple regular conjugacy classes of $G(\mathbb{C})$. Choose for each $i$ an irreducible $G(\mathbb{C})$-equivariant local system $\mathcal{L}_{i}$ on $\mathcal{C}_{i}$ and denote by $\mathcal{L}_{\mathcal{C}}$ the resulting local system on $\mathcal{M}_{\mathcal{C}}(\mathbb{C})$.
Choose a generic $k$-tuple of $F$-stable semisimple regular conjugacy classes of $G(\overline{\mathbb{F}}_{q})$ whose $i$-th coordinate is degenerate (i.e. equals to $\mathcal{C}_{-1}$) if and only if $\mathcal{C}_{i}$ is degenerate. To ease the notation we will denote this $k$-tuple again by $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$. On each conjugacy class we choose the same local system as in the complex case and denote it again by $\mathcal{L}_{i}$. We then denote by ${\bf Y}_{i}:G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}\simeq\mathbb{C}$ the characteristic function ${\bf Y}_{(\mathcal{C}_{i},\mathcal{L}_{i})}$ and by ${\bf X}_{i}$ is the characteristic function of the character-sheaf of $G^{*}$ which corresponds to $(\mathcal{C}_{i},\mathcal{L}_{i})$ under the bijection $\Phi$.
Our main theorem is the following one (see §5.1):
Theorem 1.2.1.
We have
$H_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{L}_{\mathcal{C}};q,-1)=\left\langle{\bf Y}_{1}*\cdots*{\bf Y}_{k},1_{\{1\}}\right\rangle_{G^{F}}.$
(2) The conjectured formula for the pure part $PH_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{L}_{\mathcal{C}};q)$ equals
$q^{k-3}\left\langle{\bf X}_{1}\cdots{\bf X}_{k},1\right\rangle_{G^{*}{{}^{F}}}.$
Our result reminds us of a Fourier transform.
2 $\mathrm{GL}_{2}$-character varieties
2.1 Preliminaries on cohomology
In the following, $K$ is an algebraically closed field of $\operatorname{char}\neq 2$, which is either $\mathbb{C}$ or $\overline{\mathbb{F}}_{q}$ and $X$ an algebraic variety over $K$. We denote by $H^{*}_{c}(X)$ its compactly supported cohomology, with $\mathbb{C}$ coefficients if $K=\mathbb{C}$ and $\overline{{\mathbb{Q}}}_{\ell}$-coefficients if $K=\overline{\mathbb{F}}_{q}$.
If $K=\mathbb{C}$, each cohomology group $H^{k}_{c}(X)$ is equipped with the weight filtration $W^{i}_{\bullet}$, introduced by Deligne in [7].
If $K=\overline{\mathbb{F}}_{q}$ and if we assume given a geometric Frobenius $F:X\rightarrow X$ (or, equivalently, an $\mathbb{F}_{q}$-scheme $X_{o}$ such that $X=X_{o}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$) we have a weight filtration $W^{k}_{\bullet}H^{k}_{c}(X)$, where $W^{k}_{m}H^{k}_{c}(X)$ is the subspace on which the eigenvalues of the Frobenius $F$ are of absolute value $\leq q^{\frac{m}{2}}$.
In both cases, we define the mixed Poincaré polynomial $H_{c}(X;q,t)\in\mathbb{Z}[\sqrt{q},t]$
$H_{c}(X;q,t)\coloneqq\sum_{i,k}{\rm dim}(W^{k}_{i}/W^{k}_{i-1})q^{i/2}t^{k}.$
Then $H_{c}(X;1,t)=\sum_{k}{\rm dim}\,H_{c}^{k}(X)\,t^{k}$ is the compactly supported Poincaré polynomial and $H_{c}(X;q,-1)$ is the so-called $E$-polynomial denoted by $E(X;q)$. In our cases, $H_{c}(X,q,t)$ and $E(X,q)$ will be actual polynomial in $q$, i.e. if $W^{k}_{i}/W^{k}_{i-1}\neq 0$ then $i$ is even.
We also define the pure part $PH_{c}(X;q)$ as
$PH_{c}(X;q)\coloneqq\sum_{k}\dim(W^{k}_{k}/W^{k}_{k-1})q^{k/2}.$
2.1.1 Polyomial count varieties
Given a variety $X/_{\overline{\mathbb{F}}_{q}}$ with Frobenius $F$, we say that $X/_{\overline{\mathbb{F}}_{q}}$ has polynomial count if there exists $P_{X}(t)\in\mathbb{Z}[t]$ such that
$P_{X}(q^{m})=|X^{F^{m}}|=|X_{o}(\mathbb{F}_{q^{m}})|$
for all integer $m>0$.
In this case, we have the following [20, Theorem 2.2.6].
Theorem 2.1.1.
If $X/_{\overline{\mathbb{F}}_{q}}$ has polynomial count with counting polynomial $P_{X}(t)$, we have
$P_{X}(q)=E(X;q).$
Example 2.1.2.
The variety $X=\overline{\mathbb{F}}_{q}^{*}$, with the Frobenius $F(z)=z^{q}$ has polynomial count with counting polynomial $P_{X}(t)=t-1$. Indeed, in this case we have
$|X^{F^{m}}|=|\mathbb{F}_{q^{m}}^{*}|=q^{m}-1.$
Notice that $H_{c}(X;q,t)=qt^{2}+t$ and thus $E(X;q)=P_{X}(q)$.
For a variety $X/_{\mathbb{C}}$, we say that $X$ has polynomial count with counting polynomial $P_{X}(t)$ if there exists a finitely generated $\mathbb{Z}$-subalgebra $R\subseteq\mathbb{C}$ and an $R$-scheme $X_{R}$ such that
$X_{R}\times_{R}\mathbb{C}\cong X,$
and such that, for any ring homomorphism $\varphi:R\to\mathbb{F}_{q}$, the variety $X^{\varphi}=X_{R}\times_{R}\overline{\mathbb{F}}_{q}$ has polynomial count with counting polynomial $P_{X}(t)$.
We have the following result.
Theorem 2.1.3.
[10, Appendix by Katz] If $X/_{\mathbb{C}}$ has polynomial count with counting polynomial $P_{X}(t)$, then
$E(X;q)=P_{X}(q).$ (2.1.1)
Katz’s formulation of the above theorem gives also information on the Hodge filtration. If we do not bother about the Hodge filtration (as it is the case in this paper) we can prove the above theorem in a simple way as follows.
Fix an isomorphism $\mathbb{C}\cong\overline{{\mathbb{Q}}}_{\ell}$ and identify $H^{*}_{c}(X,\mathbb{C})\cong H^{*}_{c}(X,\overline{{\mathbb{Q}}}_{\ell})$ through this isomorphism. Recall that, if $\operatorname{char}(\overline{\mathbb{F}}_{q})$ is sufficiently large, we have a natural isomorphism
$H^{*}_{c}(X^{\varphi})\cong H^{*}_{c}(X).$ (2.1.2)
Indeed, let $f:X_{R}\to\operatorname{Spec}(R)$ be the structural morphism. The complex $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constructible, see for instance [5, Chapter 2]. In particular, there exists a non-empty open $U\subseteq\operatorname{Spec}(R)$ such that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant. Denote by $\xi:\operatorname{Spec}(\mathbb{C})\to\operatorname{Spec}(R)$ the (geometric) generic point of $\operatorname{Spec}(R)$ coming from the embedding $R\subseteq\mathbb{C}$ and, for any $\varphi:R\to\mathbb{F}_{q}$, denote by $\xi_{\varphi}:\operatorname{Spec}(\overline{\mathbb{F}}_{q})\to\operatorname{Spec}(R)$ the corresponding geometric point.
If $\operatorname{Im}(\xi_{\varphi})\in U$, from the fact that $Rf_{!}\overline{{\mathbb{Q}}}_{\ell}$ is constant and from proper base change, we have the following chain of isomorphisms:
$H^{*}_{c}(X^{\varphi})\cong(Rf_{!}\overline{{\mathbb{Q}}}_{\ell})_{\xi_{\varphi}}\cong(Rf_{!}\overline{{\mathbb{Q}}}_{\ell})_{\xi}\cong H^{*}_{c}(X).$ (2.1.3)
The results of [6, Theorem 14] show that the isomorphism (2.1.2) preserves the weight filtration on both sides. In particular, we have an equality
$H_{c}(X;q,t)=H_{c}(X^{\varphi};q,t).$ (2.1.4)
We deduce thus Formula (2.1.1) from Theorem 2.1.1 and formula (2.1.4).
2.2 Review on cohomology of $\mathrm{GL}_{n}$-character varieties
To alleviate the notation we will put $\operatorname{GL}_{n}=\operatorname{GL}_{n}(K)$. If $K=\overline{\mathbb{F}}_{q}$, we denote by $F:\operatorname{GL}_{n}\to\operatorname{GL}_{n}$ the standard Frobenius $F((a_{i,j}))=(a_{i,j}^{q})$.
Let $C$ be a $k$-tuple $C=(C_{1},\dots,C_{k})$ of semisimple conjugacy classes of $\operatorname{GL}_{n}$. If $K=\overline{\mathbb{F}}_{q}$, we assume that each $C_{i}$ is $F$-stable and split, i.e. its eigenvalues belong to $\mathbb{F}_{q}^{*}$.
We define the corresponding $\operatorname{GL}_{n}$-character stack $\mathcal{M}_{C}$ as the quotient stack
$\mathcal{M}_{C}=\left[\{(x_{1},\dots,x_{k})\in C_{1}\times\cdots\times C_{k}\,|\,x_{1}\cdots x_{k}=1\}/\operatorname{PGL}_{n}\right]$
for the diagonal conjugation action of $\operatorname{PGL}_{n}$.
Following [11, Definition 2.1.1], we have the following definition of a generic $k$-tuple $C$.
Definition 2.2.1.
The $k$-tuple $C$ is generic if
$\prod_{i=1}^{k}\det(C_{i})=1,$ (2.2.1)
and for any integer $1\leq r<n$, if we select $r$ eigenvalues of $C_{i}$ (for each $i=1,\dots,k$), the product of the $nr$ selected eigenvalues is different from $1$.
For each $i$ let $\mu^{i}$ be the partition of $n$ whose parts are given by the multiplicities of the eigenvalues of $C_{i}$. We say that $C_{i}$ is of type $\mu^{i}$ and that the $k$-tuple $C$ is of type ${\bm{\mu}}=(\mu^{1},\dots,\mu^{k})$.
Remark 2.2.2.
Recall [11, Lemma 2.1.2] that for any $k$-tuple ${\bm{\mu}}=(\mu^{1},\dots,\mu^{k})$ of partitions of $n$, if $\operatorname{char}(K)$ does not divide $\gcd(\mu^{j}_{i})_{i,j}$, there always exists a generic $k$-tuple of semisimple conjugacy classes of $\operatorname{GL}_{n}$ of type $\bm{\mu}$.
Recall the following result [11, Proposition 2.1.4, Theorem 2.1.5] [12, Theorem 1.1.1].
Theorem 2.2.3.
If $C=(C_{1},\dots,C_{k})$ is generic, then, if not empty, $\mathcal{M}_{C}$ is a nonsingular irreducible affine algebraic variety of dimension
$d_{\bm{\mu}}:=(k-2)n^{2}-\sum_{i,j}(\mu^{i}_{j})^{2}+2$
where $\mu^{i}=(\mu^{i}_{1},\mu^{i}_{2},\dots)$.
In [11] the authors defined a rational function $\mathbb{H}_{\bm{\mu}}(z,w)$ which conjecturally computes the polynomial $H_{c}(\mathcal{M}_{C};q,t)$. It is defined as follows.
Consider $k$ sets ${\bm{x}}_{1},\dots,{\bm{x}}_{k}$ of infinitely many independent variables and we let $\Lambda({\bm{x}}_{1},\dots,{\bm{x}}_{k})$ be the ring of functions separately symmetric in each of the set of variables. We then put
$\Omega(z,w):=\sum_{\lambda\in\mathcal{P}}\mathcal{H}_{\lambda}(z,w)\prod_{i=1}^{k}\tilde{H}_{\lambda}({\bm{x}}_{i};z^{2},w^{2})$
where $\mathcal{P}$ is the set of all partitions, $\tilde{H}_{\lambda}({\bm{x}}_{i};z,w)$ are the modified Macdonald polynomials and
$\mathcal{H}_{\lambda}(z,w):=\prod_{s}\frac{1}{(z^{2a(s)+2}-w^{2l(s)})(z^{2a(s)}-w^{2l(s)+2})}$
where $s$ runs over the boxes of the Young diagram of $\lambda$ with $a(s)$ and $l(s)$ its arm and leg length.
Then
$\mathbb{H}_{\bm{\mu}}(z,w):=(z^{2}-1)(1-w^{2})\left\langle{\rm Log}\,\Omega(z,w),h_{\bm{\mu}}\right\rangle$
where $h_{\bm{\mu}}:=h_{\mu^{1}}({\bm{x}}_{1})\cdots h_{\mu^{k}}({\bm{x}}_{k})\in\Lambda({\bm{x}}_{1},\dots,{\bm{x}}_{k})$ is the product of complete symmetric functions and ${\rm Log}$ is the plethystic logarithm.
It is proved in [26] that these rational functions $\mathbb{H}_{\bf\mu}(z,w)$ are polynomials in $z$, $w$ with integer coefficients as it was conjectured in [11].
We have the following conjecture [11, Conjecture 1.2.1].
Conjecture 2.2.4.
We have the following.
The polynomial $H_{c}(\mathcal{M}_{C};q,t)$ depends only on the type ${\bm{\mu}}$ of $C$ and not the choice of generic eigenvalues and
$H_{c}(\mathcal{M}_{C};q,t)=(t\sqrt{q})^{d_{\bm{\mu}}}\mathbb{H}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},-t\sqrt{q}\right).$ (2.2.2)
In particular,
$PH_{c}(\mathcal{M}_{C};q)=q^{d_{\bm{\mu}}/2}\mathbb{H}_{\bm{\mu}}(0,\sqrt{q}).$ (2.2.3)
Notice that since we are working with the Riemann sphere, the sign in the second coordinate could have been omitted because in this case $\mathbb{H}_{\bm{\mu}}(z,w)$ is a function of $z^{2}$ and $w^{2}$ (which is not the case in higher genus). We decided to keep the formulas as there are in [11] for the convenience of the reader.
The Euler specialization $t\mapsto-1$ of Formula (2.2.2) is proved [11, Theorem 1.2.3] by counting points of $\mathcal{M}_{C}$ over finite fields and using Theorem 2.1.1, more precisely we have the following result.
Theorem 2.2.5.
$E(\mathcal{M}_{C};q)=q^{d_{\bm{\mu}}/2}\mathbb{H}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right).$
In particular the polynomial $E(\mathcal{M}_{C};q)$ depends only on the type ${\bm{\mu}}$ of $C$ and not the choice of generic eigenvalues.
Let us now write a more explicit formula when $n=2$.
Proposition 2.2.6.
Assume that $k\geq 1$ and ${\bm{\mu}}=((1^{2}),\dots,(1^{2}))$, then
$\mathbb{H}_{\bm{\mu}}(z,w)=\frac{(w^{2}+1)^{k}}{(z^{2}-w^{2})(1-w^{4})}+\frac{(z^{2}+1)^{k}}{(z^{4}-1)(z^{2}-w^{2})}-\frac{2^{k-1}}{(z^{2}-1)(1-w^{2})}$
Remark 2.2.7.
As mentioned earlier, we know from [26] that these rational functions are in fact polynomials in $z$ and $w$ with integer coefficients which a priori is not obvious.
Proof of Proposition 2.2.6.
The conjugacy classes of $\operatorname{GL}_{n}(\mathbb{F}_{q})$ are parametrised by the set of maps from the set of Frobenius-orbits of $\overline{\mathbb{F}}_{q}^{*}$ to the set of all partitions. The type of a such a map consists of the collection of pairs $(d,\lambda)$ with $d$ a positive integer and $\lambda$ a non-zero partition which is the image of a Frobenius-orbit of size $d$.
There are 4 types when $n=2$, i.e. $(1,1)^{2},(2,1),(1,(1^{2})),(1,(2^{1}))$. The first one is the type of a conjugacy class with two distincts eigenvalues in $\mathbb{F}_{q}^{*}$, the second one is the type of a conjugacy class whose characteristic polynomial is irreducible over $\mathbb{F}_{q}$ (i.e. it has two distinct eigenvalues $x,x^{q}$ in $\mathbb{F}_{q^{2}}^{*}$), and the last two types are the types respectively of central matrices and matrices with one Jordan block.
Then from [11, Formula (2.3.9)] we have
$\mathbb{H}_{\bm{\mu}}(z,w)=(z^{2}-1)(1-w^{2})\left\langle\sum_{\omega}C_{\omega}^{o}\mathcal{H}_{\omega}(z,w)\prod_{i=1}^{k}\tilde{H}_{\omega}({\bm{x}}_{i};z^{2},w^{2}),h_{\bm{\mu}}\right\rangle$
where the sum runs over the types of size $2$.
A direct calculation shows that
$C_{(2^{1})}^{o}=C_{(1^{2})}^{o}=1,\hskip 28.45274ptC_{(2,1)}^{o}=C_{(1,1)^{2}}^{o}=-\frac{1}{2}$
and
$\mathcal{H}_{(1^{2})}(z,w)$ $=\frac{1}{(z^{2}-w^{2})(1-w^{4})(z^{2}-1)(1-w^{2})}$
$\mathcal{H}_{(2^{1})}(z,w)$ $=\frac{1}{(z^{4}-1)(z^{2}-w^{2})(z^{2}-1)(1-w^{2})}$
$\mathcal{H}_{(1,1)^{2}}(z,w)$ $=\frac{1}{(z^{2}-1)^{2}(1-w^{2})^{2}}$
$\mathcal{H}_{(2,1)}(z,w)$ $=\frac{1}{(z^{4}-1)(1-w^{4})}$
To compute the pairing we need also to decompose the modified Macdonald symmetric functions into the basis of monomial symmetric functions $\{m_{\lambda}({\bm{x}})\}_{\lambda}$ which is orthogonal to the basis of complete symmetric functions $\{h_{\lambda}({\bm{x}})\}_{\lambda}$ :
$\tilde{H}_{(1^{2})}({\bm{x}};z,w)$ $=(w+1)m_{(1^{2})}({\bm{x}})+m_{(2^{1})}({\bm{x}})$
$\tilde{H}_{(2^{1})}({\bm{x}};z,w)$ $=(z+1)m_{(1^{2})}({\bm{x}})+m_{(2^{1})}({\bm{x}})$
$\tilde{H}_{(1,1)^{2}}({\bm{x}};z,w)$ $=m_{(2^{1})}({\bm{x}})+2m_{(1^{2})}({\bm{x}})$
$\tilde{H}_{(2,1)}({\bm{x}};z,w)$ $=m_{(2^{1})}({\bm{x}}).$
Hence in the case $n=2$ Conjecture 2.2.2 reads.
Conjecture 2.2.8.
For $k\geq 1$ and ${\bm{\mu}}=((1^{2}),\dots,(1^{2}))$ we have
$H_{c}(\mathcal{M}_{C};q,t)=\frac{(q^{2}t^{4}+qt^{2})^{k}}{q^{2}t^{6}(q^{2}t^{2}-1)(q^{2}t^{4}-1)}+\frac{t^{2k-6}(q+1)^{k}}{(q^{2}-1)(q^{2}t^{2}-1)}-\frac{(2qt^{2})^{k-1}}{qt^{4}(q-1)(qt^{2}-1)}$
2.3 Twisted mixed Poincaré polynomials
Assume that $X$ is a $K$-variety endowed with a action of a finite group $W$. If $K=\overline{\mathbb{F}}_{q}$, we assume that the action of $W$ commutes with $F$. The group action preserves the weight filtration on $H_{c}^{\bullet}(X)$ and for $w\in W$, we define the $w$-twisted mixed Poincaré polynomial as
$H_{c}^{w}(X;q,t):=\sum_{i,k}{\rm Tr}\left(w\,|\,W^{k}_{i}/W^{k}_{i-1}\right)\,q^{i/2}t^{k}.$
If $K=\overline{\mathbb{F}}_{q},$ we say that the pair $(X/_{\overline{\mathbb{F}}_{q}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$ in $\mathbb{Z}[t]$ if
$|X^{wF^{r}}|=P_{w}(q^{r}).$
for any integer $r\geq 1$.
We have the following twisted analogue of Theorem 2.1.1.
Theorem 2.3.1.
If $(X/_{\overline{\mathbb{F}}_{q}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$ for any $w\in W$ we have
$E^{w}(X;q)=P_{w}(q).$ (2.3.1)
where $E^{w}(X;q)\coloneqq H^{w}_{c}(X;q,-1)$.
The proof of this Theorem is very similar to that of [20, Theorem 2.8]. We give it below for completeness, see Proof 2.3.
Remark 2.3.2.
Given $(X/_{\overline{\mathbb{F}}_{q}},W)$ as above, for any $w\in W$, the map $wF:X\to X$ is a Frobenius morphism and gives thus another $\mathbb{F}_{q}$-structure on the $\overline{\mathbb{F}}_{q}$-variety $X$.
In particular, there exists an $\mathbb{F}_{q}$-scheme $X_{w}$ with an isomorphism $X_{w}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}\cong X$ such that, through this isomorphism, the geometric Frobenius $F_{w}$ of $X_{w}$ is identified with $wF$.
In general, the polynomial $P_{w}(t)$ is not the counting polynomial of $X_{w}$, i.e. it is not true that $|X_{w}(\mathbb{F}_{q^{r}})|=P_{w}(q^{r})$. Notice indeed that the LHS of the latter equality is $|X^{(wF)^{r}}|$, while the RHS is $|X^{wF^{r}}|$.
For a concrete example, consider , $X=\overline{\mathbb{F}}_{q}^{*}$ with $F(z)=z^{q}$ and $W=\bm{\mu}_{2}=\{1,\sigma\}$, with the action $\sigma\cdot x=x^{-1}$. The pair $(\overline{\mathbb{F}}_{q}^{*},\bm{\mu}_{2})$ has polynomial count with counting polynomials $P_{1}(t)=(t-1)$ and $P_{\sigma}(t)=(t+1)$. Indeed, we have $\sigma F(x)=x^{-q}$ and thus $X^{\sigma F}=\bm{\mu}_{q+1}$. Notice that, if $4$ does not divide $q+1$, i.e. if $-1$ is not a square in $\mathbb{F}_{q}^{*}$, we can consider $X_{\sigma}=\operatorname{Spec}(\mathbb{F}_{q}[s,t]/(s^{2}+t^{2}=1))$, with the isomorphism $X_{\sigma}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}\to\overline{\mathbb{F}}_{q}^{*}$ given by $(s,t)\to s+it$.
We have that
#$\#X_{\sigma}(\mathbb{F}_{q^{r}})=\begin{cases}q^{r}+1\text{ if }r\text{ is odd}\\ q^{r}-1\text{ if }r\text{ is even}\end{cases}$
If $K=\mathbb{C}$, we say that $(X/_{\mathbb{C}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$ if there exists a finitely generated $\mathbb{Z}$-subalgebra $R$ of $\mathbb{C}$, a separated finitely generated $R$-scheme $X_{R}$ equipped with a $W$-action which gives back $X$ with its $W$-action after scalar extensions from $R$ to $\mathbb{C}$, such that for any ring homomorphism $\varphi:R\to\mathbb{F}_{q}$, $(X^{\varphi},W)$ has polynomial counting with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$.
We have the following twisted version of Theorem 2.1.1.
Theorem 2.3.3.
Assume that $(X/_{\mathbb{C}},W)$ has polynomial count with (twisted) counting polynomials $\{P_{w}(t)\}_{w\in W}$. Then for any $w\in W$ we have
$E^{w}(X;q)=P_{w}(q).$ (2.3.2)
Theorem 2.3.3 above can be deduced from Theorem 2.3.1 as follows. Consider a variety $X/_{\mathbb{C}}$ and $R$ as above and $U\subseteq\operatorname{Spec}(R)$ as in section 2.1.1. Since the action of $W$ is defined over $R$, the isomorphism (2.1.3) is compatible with the $W$-action.
In particular, since the action of $W$ on $H^{*}_{c}(X)$ is defined over the rationals, through the isomorphism (2.1.2), we have
$\operatorname{Tr}\left(\,w\,|\,W_{i}^{k}H^{k}_{c}(X)/W^{k}_{i-1}H^{k}_{c}(X)\right)=\operatorname{Tr}\left(w\,|\,W_{i}^{k}H^{k}_{c}(X^{\varphi})/W^{k}_{i-1}H^{k}_{c}(X^{\varphi})\right)$
and thus
$H^{w}_{c}(X;q,t)=H^{w}_{c}(X^{\varphi};q,t)$
and thus
$E(X;q)=E(X^{\varphi};q).$
Proof of Theorem 2.3.1.
By Grothendieck’s trace formula, for any $r$, we have
$P_{w}(q^{r})=|X^{wF^{r}}|=\sum_{k}(-1)^{k}\operatorname{Tr}\left(wF^{r}\,|\,H^{k}_{c}(X,\overline{{\mathbb{Q}}}_{\ell})\right).$ (2.3.3)
Let $\lambda_{i,k,1}q^{\frac{i}{2}},\dots,\lambda_{i,k,s_{k,i}}q^{\frac{i}{2}}$ and $\alpha^{w}_{i,k,1},\dots,\alpha^{w}_{i,k,s_{k,i}}$ be the eigenvalues, counted with multiplicities, of $F$ and $w$ on $W^{k}_{i}/W^{k}_{i-1}$. Since $w$ and $F$ commute, up to reordering, we can assume that, for any $r\geq 1$,
$\operatorname{Tr}\left(wF^{r}\,|\,W^{k}_{i}/W^{k}_{i-1}\right)=\sum_{h=1}^{s_{k,i}}\alpha^{w}_{i,k,h}\lambda^{r}_{i,k,h}q^{\frac{ri}{2}}$
and thus
$\sum_{k}(-1)^{k}\operatorname{Tr}\left(wF^{r}\,|\,H^{k}_{c}(X)\right)=\sum_{i}\left(\sum_{k}(-1)^{k}\sum_{h=1}^{s_{k,i}}\alpha^{w}_{i,k,h}\lambda^{r}_{i,k,h}\right)q^{\frac{ir}{2}}.$
If $P_{w}(t)=\sum_{i}t^{i}c_{w,i}$, from eq.(2.3.3) we deduce that we have
$\sum_{k}(-1)^{k}\sum_{h=1}^{s_{k,i}}\alpha^{w}_{i,k,h}\lambda^{r}_{i,k,h}=\begin{cases}c_{w,\frac{i}{2}}\text{ if }i\text{ is even }\\ 0\text{ otherwise}\end{cases}.$ (2.3.4)
From [20, Lemma 2.9], we deduce that we have
$\sum_{k}(-1)^{k}\sum_{h=1}^{s_{k,i}}\alpha^{w}_{i,k,h}=\begin{cases}c_{w,\frac{i}{2}}\text{ if }i\text{ is even }\\ 0\text{ otherwise}\end{cases}.$
Since $\sum_{k}(-1)^{k}\sum_{h=1}^{s_{k,i}}\alpha^{w}_{i,k,h}=\sum_{k}(-1)^{k}\operatorname{Tr}(w|W^{k}_{i}/W^{k}_{i-1})$, we have the desired equality (2.3.1).
We keep the notation of the previous section with $n=2$. We will apply the above theorem to the following situation. We assume that $C=(C_{1},\dots,C_{k})$ is a generic $k$-tuple of semisimple regular conjugacy classes of $\operatorname{GL}_{2}$ where the $m$-th first conjugacy classes $C_{1},\dots,C_{m}$ (for some integer $m<k$) are the conjugacy class with eigenvalues $1$ and $-1$ and we let $H_{m}$ be the kernel of the morphism
[…]
where ${\bm{\mu}}_{2}$ is the multiplicative group $\{1,-1\}$.
The group $H_{m}$ acts on the character variety $\mathcal{M}_{C}$ by multiplication on the first $m$-th coordinates.
Put $\bm{\mu}=((1^{2}),\dots,(1^{2}))$ and fix $y=(y_{1},\dots,y_{m})\in H_{m}$. For each $i=1,\dots,m$, define
$h^{y_{i}}_{(1^{2})}({\bm{x}}_{i})=\begin{cases}h_{(1^{2})}({\bm{x}}_{i})&\text{ if }y_{i}=1,\\ p_{2}({\bm{x}}_{i})&\text{ otherwise}\end{cases}$
where $p_{2}({\bm{x}}_{i})=x_{i1}^{2}+x_{i2}^{2}+\cdots$ is the power sum symmetric function in the variable ${\bm{x}}_{i}=\{x_{i1},x_{i2},\dots\}$ and put
$\mathbb{H}^{y}_{\bm{\mu}}(z,w):=(z^{2}-1)(1-w^{2})\left\langle{\rm Log}\,\Omega(z,w),h^{y}_{\bm{\mu}}\right\rangle$
where
$h^{y}_{\bm{\mu}}:=h^{y_{1}}_{(1^{2})}({\bm{x}}_{1})\cdots h^{y_{m}}_{(1^{2})}({\bm{x}}_{m})h_{(1^{2})}({\bm{x}}_{m+1})\cdots h_{(1^{2})}({\bm{x}}_{k})\in\Lambda({\bm{x}}_{1},\dots,{\bm{x}}_{k}).$
Proposition 2.3.4.
Let $r=r(y):=\#\{i=1,\dots,m\,|\,y_{i}=-1\}>0$, then
$\mathbb{H}^{y}_{\bm{\mu}}(z,w)=\frac{(w^{2}+1)^{k-r}(1-w^{2})^{r}}{(z^{2}-w^{2})(1-w^{4})}+\frac{(1-z^{2})^{r}(z^{2}+1)^{k-r}}{(z^{4}-1)(z^{2}-w^{2})}.$
Proof.
Similar calculation as for the proof of Proposition 2.2.6, noticing that
$\langle\tilde{H}_{(1^{2})}({\bm{x}};z,w),p_{2}({\bm{x}})\rangle=\langle\tilde{H}_{(2,1)}({\bm{x}};z,w),h_{(1^{2})}({\bm{x}})\rangle=0$
which explains why we have only two terms. ∎
We make the following conjecture.
Conjecture 2.3.5.
We have
$H_{c}^{y}(\mathcal{M}_{C};q,t)=(t\sqrt{q})^{d_{\bm{\mu}}}\mathbb{H}_{\bm{\mu}}^{y}\left(\frac{1}{\sqrt{q}},-t\sqrt{q}\right).$
We prove that the above conjecture is true under the Euler specialization $t\mapsto-1$.
Theorem 2.3.6.
We have
$E^{y}(\mathcal{M}_{C};q)=q^{d_{\bm{\mu}}/2}\mathbb{H}^{y}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right).$
In particular, the polynomial $E^{y}(\mathcal{M}_{C};q)$ depends only on the type ${\bm{\mu}}$ of $C$ and not the choice of generic eigenvalues.
Proof.
From Formula 2.1.4, it is enough to show the case of $K=\overline{\mathbb{F}}_{q}$. Take thus a finite field $\mathbb{F}_{q}$ of characteristic $\neq 2$ and consider a generic $k$-tuple $C=(C_{1},\dots,C_{k})$ of semisimple regular conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ with eigenvalues in $\mathbb{F}_{q}$ and such that the first $m$ conjugacy classes are the conjugacy class with eigenvalues $1,-1$.
Then group $H_{m}$ acts on $\mathcal{M}_{C}$ by multiplication coordinate by coordinate on the first $m$-th coordinates.
By Theorem 2.3.1 we need to show that
#$\#(\mathcal{M}_{C})^{yF}=q^{d_{\bm{\mu}}/2}\mathbb{H}^{y}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right)$
where $F$ is the standard Frobenius that raises matrix coefficients to their $q$-th power.
Fix $i=1,\dots,m$, let $\alpha_{i}\in\overline{\mathbb{F}}_{q}^{*}$ be defined as $\alpha_{i}=1$ if $y_{i}=1$ and as an element of $\mathbb{F}_{q^{2}}^{*}$ such that $\alpha_{i}^{q}=-\alpha_{i}$.
We then let ${O}_{i}$ be the conjugacy class of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ with eigenvalues $\alpha_{i},-\alpha_{i}$ (it is the conjugacy class ${C}_{i}$ when $y_{i}=1$).
Notice that the standard Frobenius $F$ leaves stable the conjugacy classes ${O}_{i}$ and that we have the following commutative diagram
[…]
where $f_{i}$ is the multiplication by the scalar $\alpha_{i}$.
Therefore via $\prod_{i}f_{i}$, the pair $(\mathcal{M}_{C},yF)$ can be identified with $(\mathcal{M}_{O},F)$ where ${O}$ is the $k+1$-tuple of conjugacy classes
${O}=\left({O}_{1},\dots,{O}_{k},\{(\alpha_{1}\cdots\alpha_{m})^{-1}I_{2}\}\right)$
where ${O}_{i}={C}_{i}$ for $i>m$.
Notice that $\alpha_{1}\cdots\alpha_{m}\in\mathbb{F}_{q}^{\times}$ as there is an even number of non-trivial $y_{i}$.
We thus have
#$\#(\mathcal{M}_{C})^{yF}=\#(\mathcal{M}_{O})^{F}.$
Now for an $F$-stable conjugacy class ${O}$ of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ denote by $1_{\bm{O}}:\operatorname{GL}_{2}(\mathbb{F}_{q})\rightarrow\mathbb{C}^{*}$ the function that takes the value $1$ at ${\bm{O}}^{F}$ and 00 elsewhere.
Then
#$\#(\mathcal{M}_{O})^{F}=\frac{1}{|\operatorname{PGL}_{n}(\mathbb{F}_{q})|}\left(1_{{O}_{1}}*\cdots*1_{{O}_{k}}*1_{(\alpha_{1}\cdots\alpha_{k})^{-1}I_{2}}\right)(1)$
It follows from [18, Theorem 4.14] that the RHS equals to
$q^{d_{\bm{\mu}}/2}\mathbb{H}^{y}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right).$
Indeed, since the conjugacy classes ${O}_{i}$ are semisimple, the function ${\bm{X}}_{\overline{O}_{i}}$ in [18, Theorem 4.14] is our function $1_{{\bm{O}}_{i}}$. In loc. cit. we introduced the notion of type of a conjugacy class of $\operatorname{GL}_{n}(\mathbb{F}_{q})$ (see [18, §4.1.2]) which we could avoid here as we are working with semisimple conjugacy classes of $\operatorname{GL}_{2}$. The type of a split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ is $\omega=(1,1)(1,1)$ and the symmetric function $s_{\omega}$ (see [18, Remark 4.1]) is our symmetric function $h_{(1^{2})}$. The type of a non-split semisimple regular conjugacy class of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ like our conjugacy class ${O}_{i}$ with $y_{i}=-1$ (it has eigenvalues in $\mathbb{F}_{q^{2}}-\mathbb{F}_{q}$) is $\omega=(2,1)$ and $s_{\omega}$ is our symmetric function $p_{2}$. Finally in the formula for $\mathbb{H}_{\bm{\omega}}(z,w)$ in [18, Above Lemma 4.2], there is a sign which in our case is always $1$ as we have an even number of twisted conjugacy classes. ∎
3 $\mathrm{PGL}_{2}$-character varieties
3.1 Local systems on semisimple conjugacy classes of $\mathrm{PGL}_{2}$
In this section we work over an algebraically closed field $K$ of characteristic $\neq 2$ (for us $K$ will be either $\mathbb{C}$ or $\overline{\mathbb{F}}_{q}$). If $K=\overline{\mathbb{F}}_{q}$ we will work with $\ell$-adic sheaves and if $K=\mathbb{C}$ we can work either with $\ell$-adic sheaves or complex sheaves for the analytic topology. We thus let $\kappa$ be the field $\mathbb{C}$ or $\overline{{\mathbb{Q}}}_{\ell}$ depending on the context and use the same letter to denote the constant sheaf.
We denote by
$p:\operatorname{GL}_{2}(K)\to\operatorname{PGL}_{2}(K)$
the canonical projection map. Let $T\subseteq\operatorname{GL}_{2}(K)$ and $\mathcal{T}\subseteq\operatorname{PGL}_{2}(K)$ be the maximal tori of diagonal matrices.
Identify $W_{\operatorname{GL}_{2}}(T)=\mathbb{Z}/2\mathbb{Z}$ and $W_{\operatorname{PGL}_{2}}(\mathcal{T})=\mathbb{Z}/2\mathbb{Z}$ in the classical way, through permutation matrices. Notice that $p(W_{\operatorname{GL}_{2}}(T))=W_{\operatorname{PGL}_{2}}(\mathcal{T})$.
We now describe the irreducible $\operatorname{PGL}_{2}$-equivariant local systems on the regular semisimple conjugacy classes of $\operatorname{PGL}_{2}(K)$.
For $x\in K^{*}\setminus{1}$, let $g_{x}$ be the matrix
$g_{x}=\begin{pmatrix}1&0\\ 0&x\end{pmatrix}.$
The centraliser $C_{\operatorname{GL}_{2}}(g_{x})$ of $g_{x}$ in $\operatorname{GL}_{2}(K)$ is $T$ and so in particular it is connected. Therefore the only irreducible $\operatorname{GL}_{2}(K)$-equivariant local system on the $\operatorname{GL}_{2}(K)$-conjugacy class $C_{x}$ of $g_{x}$ is the constant sheaf $\kappa$.
$\bullet$ If $x\neq-1$, then $C_{\operatorname{PGL}_{2}(K)}(p(g_{x}))=\mathcal{T}$ is connected and so the only irreducible $\operatorname{PGL}_{2}(K)$-equivariant local system on the $\operatorname{PGL}_{2}(K)$-conjugacy class $\mathcal{C}_{x}$ of $p(g_{x})$ is the constant sheaf. Notice that $p$ restricts to an isomorphism
$p:C_{x}\to\mathcal{C}_{x}.$
$\bullet$ If $x=-1$, then
$C_{\operatorname{PGL}_{2}(K)}(p(g_{-1}))=\mathcal{T}\rtimes W_{\operatorname{PGL}_{2}}(\mathcal{T})=\mathcal{T}\rtimes\mathbb{Z}/2\mathbb{Z}$
has two connected components and so there are exactly two (up to isomorphism) irreducible $\operatorname{PGL}_{2}(K)$-equivariant local systems on the $\operatorname{PGL}_{2}(K)$-conjugacy class $\mathcal{C}_{-1}$ of $p(g_{-1})$. The non-trivial local system $\mathcal{L}$ is obtained from the $2$-covering
$p:C_{-1}\to\mathcal{C}_{-1}$
as
$p_{*}(\kappa)\cong\kappa\oplus\mathcal{L}.$
The corresponding non-trivial covering automorphism $\sigma:C_{-1}\to C_{-1}$ is given by $\sigma(x)=-x$.
The local system $\mathcal{L}$ has thus the following explicit description. For an open $U\subseteq\mathcal{C}_{-1}$, we have
$\mathcal{L}(U)=\{f\in\kappa(p^{-1}(U))\ |\ f\circ\sigma=-f\}.$ (3.1.1)
We will call the conjugacy classes $\mathcal{C}_{x}$, with $x\neq-1$, non-degenerate and $\mathcal{C}_{-1}$ degenerate.
3.2 Geometry of $\mathrm{PGL}_{2}$-character stacks and varieties
Let $p^{k}:\operatorname{GL}_{2}^{k}\to\operatorname{PGL}_{n}^{k}$ be the $k$-th product of the projection map. Fix a $k$-tuple $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ of regular semisimple conjugacy classes of $\operatorname{PGL}_{2}$ and consider a $k$-tuple $C=(C_{1},\dots,C_{k})$ of conjugacy classes of $\operatorname{GL}_{2}$ such that $p(C_{i})=\mathcal{C}_{i}$ for each $i=1,\dots,k$. Fix a square root
$\lambda_{C}=\sqrt{\prod_{i=1}^{k}\det(C_{i})}$
and consider the following affine algebraic varieties
$X_{\mathcal{C}}\coloneqq\left\{(x_{1},\dots,x_{k})\in\mathcal{C}_{1}\times\cdots\times\mathcal{C}_{k}\ |\ x_{1}\cdots x_{k}=1\right\}$
$X^{+}_{C}\coloneqq\{(x_{1},\dots,x_{k})\in C_{1}\times\cdots\times C_{k}\ |\ x_{1}\cdots x_{k}=\lambda_{C}I_{2}\}$
$X^{-}_{C}\coloneqq\{(x_{1},\dots,x_{k})\in C_{1}\times\cdots\times C_{k}\ |\ x_{1}\cdots x_{k}=-\lambda_{C}I_{2}\}.$
We have a decomposition
$(p^{k})^{-1}(X_{\mathcal{C}})=X^{+}_{\mathcal{C}}\bigsqcup X^{-}_{\mathcal{C}}$
Notice that $\operatorname{PGL}_{2}$ acts diagonally by conjugation on each of the above varieties. We consider the following GIT quotients
$M^{\pm}_{C}\coloneqq X^{\pm}_{C}/\!/\operatorname{PGL}_{2}$
$M_{\mathcal{C}}\coloneqq X_{\mathcal{C}}/\!/\operatorname{PGL}_{2}$
and the quotient stacks
$\mathcal{M}^{\pm}_{C}\coloneqq[X^{\pm}_{C}/\operatorname{PGL}_{2}].$
$\mathcal{M}_{\mathcal{C}}\coloneqq[X_{\mathcal{C}}/\operatorname{PGL}_{2}]$
We have the following Proposition.
Proposition 3.2.1.
(1) If the $\operatorname{GL}_{2}$-conjugacy classes $C_{1},\dots,C_{k}$ are all non-degenerate, we have isomorphisms:
$M_{\mathcal{C}}\cong M^{+}_{C}\bigsqcup M^{-}_{C}$ (3.2.1)
and
$\mathcal{M}_{\mathcal{C}}=\mathcal{M}^{+}_{C}\bigsqcup\mathcal{M}^{-}_{C}$ (3.2.2)
(2) Suppose that there are $m\geq 1$ degenerate conjugacy classes among $C_{1},\dots,C_{k}$. We have isomorphisms
$M_{C}^{+}\cong M_{C}^{-}$ (3.2.3)
$\mathcal{M}_{C}^{+}\cong\mathcal{M}^{-}_{C}.$ (3.2.4)
Moreover, denote by $p^{+}:\mathcal{M}_{C}^{+}\to\mathcal{M}_{\mathcal{C}}$ the map obtained by restricting $p^{k}$ to $X_{C}^{+}$. The map $p^{+}$ is a Galois covering with Galois group $H_{m}={\rm Ker}(\epsilon_{m})$ (see §2.3).
Proof.
We have a Cartesian diagram
[…]
From the description of the conjugacy classes given in section 3.1, if each the conjugacy classes $\mathcal{C}_{i}$ are all non-degenerate, $p^{k}$ is an isomorphism and
$X^{+}_{\mathcal{C}}\bigsqcup X^{-}_{\mathcal{C}}\cong X_{\mathcal{C}}$ (3.2.5)
Taking the quotient by $\operatorname{PGL}_{2}$, we obtain the isomorphisms (3.2.1) and (3.2.2).
In the rest of proof, we assume that $\mathcal{C}_{1},\dots,\mathcal{C}_{m}$ are the degenerate classes.
From the description of the conjugacy classes given in section 3.1, if there are $m$ degenerate conjugacy classes, we see that $p^{k}$ is a Galois covering with group $({\bm{\mu}}_{2})^{m}$. The $({\bm{\mu}}_{2})^{m}$-action is defined as follows.
For $y=(y_{1},\dots,y_{m})\in({\bm{\mu}}_{2})^{m}$ and $(x_{1},\dots,x_{k})\in C_{1}\times\cdots\times C_{k}$, we have
$y\cdot(x_{1},\dots,x_{k})=(y_{1}x_{1},y_{2}x_{2},\dots,y_{m}x_{m},x_{m+1},\dots,x_{k}).$
By pullback, we see that the map
$X_{C}^{+}\sqcup X_{C}^{-}\to X_{\mathcal{C}}$
is a $({\bm{\mu}}_{2})^{m}$-Galois covering too. Consider the element $\tau=(-1,1,\dots,1)\in({\bm{\mu}}_{2})^{m}$ and let $H_{m}^{\prime}$ be the subgroup of $({\bm{\mu}}_{2})^{m}$ generated by $\tau$. Notice that
$({\bm{\mu}}_{2})^{m}\cong H_{m}\oplus H_{m}^{\prime}$ (3.2.6)
The automorphism associated to $\tau$ restricts to an isomorphism
$\tau:X_{C}^{+}\to X_{C}^{-}$
$(x_{1},\dots,x_{k})\to(-x_{1},\dots,x_{k}).$
It is $\operatorname{PGL}_{2}$-equivariant from which we get the isomorphisms (3.2.3) and (3.2.4).
From the splitting (3.2.6), we see that
$X_{C}^{+}\to X_{\mathcal{C}}$
is a Galois covering with Galois group $H_{m}$. Since the Galois covering structure is $\operatorname{PGL}_{2}$-equivariant, we deduce that the quotient map $p^{+}$ is also a Galois covering with Galois group $H_{m}$.
Remark 3.2.2.
Consider the case in which $m=1$, i.e. there is exactly one degenerate class among $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$. From Proposition 3.2.1, we see that we have an isomorphism
$\mathcal{M}^{+}_{C}\cong\mathcal{M}_{\mathcal{C}}.$
Consider the $k+1$-tuples of conjugacy classes
$C^{+}=(C_{1},\dots,C_{k},\lambda_{C}^{-1}I_{2}),\,\,\,\,C^{-}=(C_{1},\dots,C_{k},-\lambda_{C}^{-1}I_{2})$
of $\operatorname{GL}_{2}$.
Definition 3.2.3.
The $k$-tuple $\mathcal{C}$ of conjugacy classes of $\operatorname{PGL}_{2}$ is said to be generic if the $(k+1)$-tuples $C^{+},C^{-}$ are both generic in the sense of §2.2.
Remark 3.2.4.
Notice that, if at least one of the conjugacy class $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ is degenerate (i.e. if $m\geq 1$), then $C^{+}$ is generic if and only if $C^{-}$ is generic.
Remark 3.2.5.
In [13, Definition 9], the authors give a definition of a generic $k$-tuple of conjugacy classes for any reductive group $G$. It is not hard to see that their definition agrees with our definition in the case of $\operatorname{PGL}_{2}$.
From [11, Theorem 2.1.5] and Proposition 3.2.1 we deduce the following Theorem.
Theorem 3.2.6.
If $\mathcal{C}$ is generic, the $\operatorname{PGL}_{2}$-character stack $\mathcal{M}_{\mathcal{C}}$ is a smooth algebraic variety (i.e. the canonical morphism $\mathcal{M}_{\mathcal{C}}\rightarrow M_{\mathcal{C}}$ is an isomorphism of stacks) of dimension $62k-$ 6.
3.3 Local systems on $\mathrm{PGL}_{2}$-character stacks
We keep the notation of the previous section and we fix a non-negative integer $m<k$ and we assume that the first $m$ conjugacy classes of $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ are degenerate and the last $k-m$ conjugacy classes are non-degenerate. From Proposition 3.2.1, the map $p^{+}:\mathcal{M}^{+}_{C}\to\mathcal{M}_{\mathcal{C}}$ is a Galois covering with Galois group $H_{m}$.
We have therefore a decomposition
$(p^{+})_{*}\kappa\cong\bigoplus_{\chi\in\widehat{H_{m}}}\mathcal{F}_{\chi},$ (3.3.1)
where $\widehat{H_{m}}$ denotes the set of irreducible characters of $H_{m}$ and where each $\mathcal{F}_{\chi}$ is an irreducible local system. We have for any open subset $U$ of $\mathcal{M}_{\mathcal{C}}$
$\mathcal{F}_{\chi}(U)=\{f\in\kappa((p^{+})^{-1}(U))\ |\ y^{*}(f)=\chi(h)f\ \ \text{ for each }y\in H_{m}\}$ (3.3.2)
where, for $y\in H_{m}$, we denote by $y:\mathcal{M}^{+}_{C}\to\mathcal{M}^{+}_{C}$ the corresponding covering automorphism.
Let us now describes the set of irreducible characters of $H_{m}$. For each subset $A\subseteq\{1,\dots,m\}$, define $\chi_{A}\in\widehat{({\bm{\mu}}_{2})^{m}}$ as
$\chi_{A}:({\bm{\mu}}_{2})^{m}\to\mathbb{C}^{\times}$
$(y_{1},\dots,y_{m})\mapsto\prod_{j\in A}y_{j}.$
Example 3.3.1.
With the notation just introduced, we have that $\chi_{\varnothing}=1$ and $\chi_{\{1,\dots,m\}}=\epsilon_{m}$.
Notice that $\widehat{({\bm{\mu}}_{2})^{m}}=\{\chi_{A}\}_{A\subseteq\{1,\dots,m\}}$ and we have the following Proposition.
Proposition 3.3.2.
The restriction map
$\operatorname{Res}_{H_{m}}:\widehat{({\bm{\mu}}_{2})^{m}}\to\widehat{H_{m}}$
is surjective and we have $\operatorname{Res}_{H_{m}}(\chi_{A})=\operatorname{Res}_{H_{m}}(\chi_{B})$ if and only if $\chi_{A}=\chi_{B}$ or $\chi_{A}=\epsilon_{m}\chi_{B}$, i.e. if and only if $A=B$ or $B=A^{c}$.
For $B\subseteq\{2,\dots,m\}$ we denote $\operatorname{Res}_{H_{m}}(\chi_{B\cup\{1\}})$ by $\overline{\chi_{B}}$. We have therefore the following bijection
$\widehat{H_{m}}\longleftrightarrow\{\overline{\chi_{B}}\ |\ B\subseteq\{2,\dots,m\}\}.$
The decomposition of (3.3.1) can therefore be rewritten as
$(p^{+})_{*}\kappa\cong\bigoplus_{B\subseteq\{2,\dots,m\}}\mathcal{F}_{\overline{\chi_{B}}}.$ (3.3.3)
We now give a concrete description of the local systems $\mathcal{F}_{\overline{\chi_{B}}}$.
Consider the closed embedding
$i_{\mathcal{C}}:X_{\mathcal{C}}\to\mathcal{C}_{1}\times\cdots\times\mathcal{C}_{k}.$
For any $B\subseteq\{2,\dots,m\}$, put
$\mathcal{L}_{B}\coloneqq i_{\mathcal{C}}^{*}\left(\mathcal{E}_{1}\boxtimes\mathcal{E}_{2}\boxtimes\cdots\boxtimes\mathcal{E}_{k}\right)$
where
$\mathcal{E}_{i}=\begin{cases}\mathcal{L}&\text{ if }i\in\{1\}\cup B,\\ \kappa&\text{ otherwise}.\end{cases}$
Notice that $\mathcal{L}_{B}$ is $\operatorname{PGL}_{2}$-equivariant and defines therefore a local system $\mathcal{F}_{B}$ on $\mathcal{M}_{\mathcal{C}}$. From (3.1.1) and (3.3.2), we see that for any $B\subseteq\{2,\dots,m\}$, we have an equality
$\mathcal{F}_{B}=\mathcal{F}_{\overline{\chi_{B}}}.$ (3.3.4)
The decomposition (3.3.3) can therefore be rewritten as
$(p^{+})_{*}\kappa\cong\bigoplus_{B\subseteq\{2,\dots,m\}}\mathcal{F}_{B}.$ (3.3.5)
Remark 3.3.3.
Notice that if $m=1$, then $B$ is necessarily the empty set and
$\mathcal{L}_{B}=(i_{\mathcal{C}})^{*}(\mathcal{L}\boxtimes\kappa\boxtimes\cdots\boxtimes\kappa)$
is the constant sheaf.
3.4 Cohomology of generic $\mathrm{PGL}_{2}$-character varieties
In this section, we give our main results concerning the cohomology of $\operatorname{PGL}_{2}$-character stacks.
If $K=\overline{\mathbb{F}}_{q}$, we pick a generic $k$-tuple $\mathcal{C}$ of regular semisimple conjugacy classes of $\operatorname{PGL}_{2}(\overline{\mathbb{F}}_{q})$ which are $F$-stable and split, i.e. each $\mathcal{C}_{i}$ is the conjugacy class of $p(g_{x_{i}})$ with $x_{i}\in\mathbb{F}_{q}^{*}$.
Notice that, for such a $\mathcal{C}$, we can find a $k$-tuple $C$ of conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ which are $F$-stable and split semisimple and such that $p(C_{i})=\mathcal{C}_{i}$. For instance we can take $C_{i}$ to be the conjugacy class of $g_{x_{i}}$. In this case, we assume in what follows that $\lambda_{C}\in\mathbb{F}_{q}^{*}$, or, equivalently, that
$\prod_{i}\det(C_{i})\in(\mathbb{F}_{q}^{*})^{2},$
where $(\mathbb{F}_{q}^{*})^{2}\subseteq\mathbb{F}_{q}^{*}$ is the subgroup of squares.
Notice that, under these assumptions, the constructions of section 3.2 and section 3.3 are all compatible with $F$.
Remark 3.4.1.
If $\lambda_{C}\not\in\mathbb{F}_{q}^{*}$, then $|\mathcal{M}_{\mathcal{C}}^{F}|=0$, see section 5.6 for more details.
3.4.1 The non-degenerate case
Let us first discuss the simpler case where the conjugacy classes $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ are all non-degenerate.
From Proposition 3.2.1 and Theorem 2.2.5, we deduce the following Theorem.
Theorem 3.4.2.
If $\mathcal{C}$ is a generic $k$-tuple of non-degenerate regular semisimple conjugacy classes of $\operatorname{PGL}_{2}$, we have
$E(\mathcal{M}_{\mathcal{C}};q)=2q^{k-3}\mathbb{H}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right)$ (3.4.1)
The following conjecture is a consequence of the conjectural formula (2.2.2) and Proposition 3.2.1.
Conjecture 3.4.3.
If $\mathcal{C}$ is a generic $k$-tuple of non-degenerate regular semisimple conjugacy classes of $\operatorname{PGL}_{2}$, we have
$H_{c}(\mathcal{M}_{\mathcal{C}};q,t)=2(qt^{2})^{k-3}\mathbb{H}_{\bm{\mu}}\left(\frac{1}{\sqrt{q}},-t\sqrt{q}\right)$ (3.4.2)
In particular, the pure part is given by the formula
$PH_{c}(\mathcal{M}_{\mathcal{C}};q)=2q^{k-3}\mathbb{H}_{\bm{\mu}}(0,\sqrt{q}).$
3.4.2 The degenerate case
Fix now $1\leq m<k$ and consider the case where the first $m$ conjugacy classes of $\mathcal{C}_{1},\dots,\mathcal{C}_{k}$ are degenerate.
Taking cohomology in the decomposition (3.3.5) we have the following isomorphism of $H_{m}$-representations:
$H_{c}^{\bullet}(\mathcal{M}^{+}_{C})\cong\bigoplus_{B\subseteq\{2,\dots,m\}}H_{c}^{\bullet}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B})\otimes\overline{\chi_{B}}.$ (3.4.3)
i.e. $H^{\bullet}_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B})$ is the subspace of $H^{\bullet}_{c}(\mathcal{M}_{C}^{+})$ on which the group $H_{m}$ acts by the character $\overline{\chi_{B}}$.
The decomposition (3.4.3) respects the weight filtrations on both sides and so we have an equality of twisted mixed Poincaré series (see §2.3), i.e. for any $y\in H_{m}$,
$H_{c}^{y}(\mathcal{M}^{+}_{C};q,t)=\sum_{B\subseteq\{2,\dots,m\}}H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q,t)\,\overline{\chi_{B}}(y),$ (3.4.4)
where for a local system $\mathcal{L}$ on $X$ we put
$H_{c}(X,\mathcal{L};q,t):=\sum_{i,k}{\rm dim}(W^{k}_{i}/W^{k}_{i-1})q^{i/2}t^{k}$
with $W^{k}_{\bullet}$ the weight filtration on $H^{k}_{c}(X,\mathcal{L})$.
Given the orthogonality relations in the character ring of $H_{m}$, we can "invert" the isomorphism (3.4.3) and obtain that for each $B\subseteq\{2,\dots,m\}$, we have
$H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q,t)=\frac{1}{|H_{m}|}\sum_{y\in H_{m}}H^{y}_{c}(\mathcal{M}^{+}_{C};q,t)\,\overline{\chi_{B}}(y)$ (3.4.5)
The equation (3.4.5) allows to express the cohomology of the $\operatorname{PGL}_{2}$-character stack $\mathcal{M}_{\mathcal{C}}$ with coefficient in the local system $\mathcal{F}_{B}$ just in terms of the cohomology of $\mathcal{M}^{+}_{C}$ and its $H_{m}$-action.
For any $m_{1},m_{2},r\in\mathbb{N}$, denote by $C_{m_{1},m_{2},r}$ the coefficient of $y^{r}x^{m_{1}+m_{2}-r}$ in the product $(x-y)^{m_{1}}(x+y)^{m_{2}}$, and define
$\mathbb{A}_{r}(z,w):=\begin{cases}\frac{(w^{2}+1)^{k-r}(1-w^{2})^{r}}{(z^{2}-w^{2})(1-w^{4})}+\frac{(1-z^{2})^{r}(z^{2}+1)^{k-r}}{(z^{4}-1)(z^{2}-w^{2})}&\text{ if }0<r\leq m,\\ &\\ \frac{(w^{2}+1)^{k}}{(z^{2}-w^{2})(1-w^{4})}+\frac{(z^{2}+1)^{k}}{(z^{4}-1)(z^{2}-w^{2})}-\frac{2^{k-1}}{(z^{2}-1)(1-w^{2})}&\text{ if }r=0.\end{cases}$
Remark 3.4.4.
Notice that $\mathbb{A}_{0}(z,w)=\mathbb{H}_{\bm{\mu}}(z,w)$ for ${\bm{\mu}}=((1^{2}),\dots,(1^{2}))$ (see Proposition 2.2.6) and for $r>0$, $\mathbb{A}_{r}(z,w)=\mathbb{H}_{\bm{\mu}}^{y}(z,w)$ with $y\in H_{m}$ such that $r(y)=r$ (see Proposition 2.3.4).
We have the following results concerning the cohomology of the local systems $\mathcal{F}_{B}$.
Theorem 3.4.5.
For any $B\subseteq\{2,\dots,m\}$, we have
$E(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q):$ $=H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q,-1)$
$=\frac{q^{k-3}}{2^{m-1}}\sum_{\begin{subarray}{c}r=0\\ r\text{ even}\end{subarray}}^{m}C_{|B|+1,m-|B|-1,r}\,\,\mathbb{A}_{r}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right).$
Proof.
From Formula (3.4.5) and Theorem 2.3.6 we have
$E(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q)$ $=\frac{1}{|H_{m}|}\sum_{y\in H_{m}}E^{y}(\mathcal{M}^{+}_{C};q)\,\overline{\chi_{B}}(y)$
$=\frac{q^{d_{\bm{\mu}}/2}}{|H_{m}|}\sum_{y\in H_{m}}\mathbb{H}^{y}_{{\bm{\mu}}}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right)\,\overline{\chi_{B}}(y)$
First of all notice that $d_{\bm{\mu}}=2k-6$ when ${\bm{\mu}}=((1^{2}),\dots,(1^{2}))$ and from Proposition 2.3.4 we see that $\mathbb{H}^{y}_{\bm{\mu}}(z,w)=\mathbb{A}_{r(y)}(z,w)$ depends only on the number $r(y)$ of $i=1,\dots,m$ such that $y_{i}=-1$ (notice that $r(y)$ must be even to have $y\in H_{m}$). Fix an even number $r$ and $y\in H_{m}$ such that $r(y)=r$. The value $\overline{\chi_{B}}(y)$ depends on the parity of #$\#\{i\in B\cup\{1\}\,|\,y_{i}=-1\}$, it equals $1$ if this number is even and it equals $-1$ is this number is odd.
For any even $r\in\{0,\dots,m\}$, we thus have
$\sum_{\begin{subarray}{c}y\in H_{m}\\ r(y)=r\end{subarray}}\overline{\chi_{B}}(y)\,\mathbb{H}^{y}_{\bm{\mu}}(z,w)$ $=\mathbb{A}_{r}(z,w)\sum_{\begin{subarray}{c}y\in H_{m}\\ r(y)=r\end{subarray}}(-1)^{\#\{i\in B\cup\{1\}\,|\,y_{i}=-1\}}$
$=\mathbb{A}_{r}(z,w)\sum_{\begin{subarray}{c}A\subseteq\{1,\dots,m\}\\ |A|=r\end{subarray}}(-1)^{|A\cap(B\cup\{1\})|}$
$=C_{|B|+1,m-|B|-1,r}\,\mathbb{A}_{r}(z,w).$
The following conjecture is a consequence of Conjecture 2.3.5 and Formula (3.4.5).
Conjecture 3.4.6.
For any $B\subseteq\{2,\dots,m\}$, we have
$H_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q,t)=\frac{(qt^{2})^{k-3}}{2^{m-1}}\sum_{\begin{subarray}{c}r=0\\ r\text{ even}\end{subarray}}^{m}C_{|B|+1,m-|B|-1,r}\,\,\mathbb{A}_{r}\left(\frac{1}{\sqrt{q}},-t\sqrt{q}\right).$ (3.4.6)
In particular, the Poincaré polynomial of the pure part is given by
$PH_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B};q)=\frac{(qt^{2})^{k-3}}{2^{m-1}}\sum_{\begin{subarray}{c}r=0\\ r\text{ even}\end{subarray}}^{m}C_{|B|+1,m-|B|-1,r}\,\,\mathbb{A}_{r}\left(0,\sqrt{q}\right).$ (3.4.7)
4 Review on Deligne-Lusztig theory and Lusztig’s character-sheaves theory for ${\rm GL}_{2}$ and ${\rm SL}_{2}$
In this section we review the construction of the complex character table of $\operatorname{GL}_{2}(\mathbb{F}_{q})$, $\operatorname{SL}_{2}(\mathbb{F}_{q})$ as well as the construction of the character-sheaves by Lusztig [21][22][25].
We assume that $q$ is odd.
On $G=\operatorname{GL}_{n}(\overline{\mathbb{F}}_{q}),\,\operatorname{SL}_{n}(\overline{\mathbb{F}}_{q})$ we will only consider the standard Frobenius $F:G\rightarrow G$ that raises matrix coefficients to their $q$-th power so that the finite group $G^{F}$ of points fixed by $F$ is exactly matrices with coefficients in the finite field $\mathbb{F}_{q}$.
As we will work with $\ell$-adic sheaves, we need to choose a prime $\ell$ which does not divide $q$.
4.1 Deligne-Lusztig theory for $\mathrm{GL}_{2}$ and $\mathrm{SL}_{2}$
Consider a connected reductive algebraic group $G$ over $\overline{\mathbb{F}}_{q}$ with a geometric Frobenius $F:G\rightarrow G$, an $F$-stable maximal torus $T$ of $G$, a Borel subgroup $B$ containing $T$ and we denote by $U$ the unipotent radical of $B$. We have $B=T\ltimes U$. For $G=\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$, an example of such a pair would be the upper triangular matrices for $B$ and the diagonal matrices for $T$ in which case both $T$ and $B$ are $F$-stable if $F$ is the standard Frobenius, but we will need also $F$-stable maximal tori which are not contained in $F$-stable Borel subgroups.
Denote by
$L:G\rightarrow G,\hskip 14.22636ptg\mapsto g^{-1}F(g)$
the Lang map (it is surjective by a theorem of Steinberg).
The affine algebraic variety $L^{-1}(U)$ is equipped with an action of $G^{F}$ by left multiplication and with an action of $T^{F}$ by right multiplication (because $U$ is a normal subgroup of $B$). These actions induce actions on the compactly supported $\ell$-adic cohomology groups $H_{c}^{i}(L^{-1}(U),\overline{{\mathbb{Q}}})$.
For a linear character $\theta$ of $T^{F}$ we define the Deligne-Lusztig virtual character $R_{T}^{G}(\theta)$ of $G^{F}$
$R_{T}^{G}(\theta)(g)=\sum_{i}(-1)^{i}\operatorname{Tr}\left(g,\,H_{c}^{i}(L^{-1}(U),\overline{{\mathbb{Q}}}_{\ell})\otimes_{\overline{{\mathbb{Q}}}_{\ell}[T^{F}]}\theta\right)$
for all $g\in G^{F}$.
These virtual characters were defined by Deligne and Lusztig [8] (see also [9]). These are very important in the classification of the irreducible characters of $G^{F}$ (analogue of Jordan decomposition of irreducible characters by Lusztig).
When $G=\operatorname{GL}_{n}$, we can can express all irreducible characters of $G^{F}$ as a linear combination of these Deligne-Lusztig characters (this is not the case for $\operatorname{SL}_{2}$ as we will see below).
When the maximal torus $T$ is contains in some $F$-stable Borel subgroup $B$ (we say that $T$ is split), then $R_{T}^{G}(\theta)$ as the following more concrete description (see [9, below Definition 11.1]):
Consider the canonical projection $\pi:B^{F}\rightarrow T^{F}$, then
$R_{T}^{G}(\theta)={\rm Ind}_{B^{F}}^{G^{F}}(\pi^{*}(\theta))$
where ${\rm Ind}_{B^{F}}^{G^{F}}$ is the classical induction in representations of finite groups. If moreover $\theta$ is the trivial character then $R_{T}^{G}(\theta)$ is the character afforded by the $G^{F}$-module ¯$\overline{{\mathbb{Q}}}_{\ell}[G^{F}/B^{F}]$ where $G^{F}$ acts by left multiplication on $G^{F}/B^{F}$.
We have the following result.
Proposition 4.1.1.
[9, Proposition 13.22] Let $T$ be an $F$-stable maximal torus of $\operatorname{GL}_{n}$ and denote by $T^{\prime}$ the maximal torus $T\cap\operatorname{SL}_{n}$ of $\operatorname{SL}_{n}$. Let $\theta$ be a linear character of $T^{F}$, then
$R_{T^{\prime}}^{\operatorname{SL}_{n}}(\theta|_{T^{\prime}{{}^{F}}})=R_{T}^{\operatorname{GL}_{n}}(\theta)|_{\operatorname{SL}_{n}^{F}}.$
When $G=\operatorname{GL}_{2}$, we have two $G^{F}$-conjugacy classes of $F$-stable maximal tori whose representatives are chosen as follows
$T_{1}=\left.\left\{\left(\begin{array}[]{cc}a&0\\ 0&b\end{array}\right)\right|\,a,b\in\overline{\mathbb{F}}_{q}^{*}\right\},\hskip 14.22636ptT_{\sigma}=\left.\Biggl{\{}\dfrac{1}{\alpha^{q}-\alpha}\begin{pmatrix}a\alpha^{q}-b\alpha&-a+b\\ (a-b)\alpha^{q+1}&-a\alpha+b\alpha^{q}\end{pmatrix}\ \right|\ a,b\in\overline{\mathbb{F}}_{q}^{*}\Biggr{\}}.$
where $\alpha$ is a fixed element of $\mathbb{F}_{q^{2}}$ which is not in $\mathbb{F}_{q}$.
Here $\sigma$ denotes the non-trivial element of the Weyl group $W=S_{2}$ of $\operatorname{GL}_{2}$ with respect to $T_{1}$.
Then $T_{1}^{F}\simeq\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$ and
$T_{\sigma}^{F}=\Biggl{\{}\dfrac{1}{\alpha^{q}-\alpha}\begin{pmatrix}a\alpha^{q}-b\alpha&-a+b\\ (a-b)\alpha^{q+1}&-a\alpha+b\alpha^{q}\end{pmatrix}\ |\ a\in\mathbb{F}_{q^{2}}^{*}\text{ and }b=a^{q}\Biggr{\}}\simeq\mathbb{F}_{q^{2}}^{*}.$
Notice that the diagram
[…]
where the vertical arrows are the conjugation $t\mapsto gtg^{-1}$ by $g=\left(\begin{array}[]{cc}1&1\\ \alpha&\alpha^{q}\end{array}\right)$, commutes since $g^{-1}F(g)=\sigma$.
A matrix in $T_{\sigma}^{F}$ is thus geometrically conjugate to a diagonal matrix with eigenvalues of the form $x$ and $x^{q}$ with $x\in\mathbb{F}_{q^{2}}^{*}$.
When $G=\operatorname{SL}_{2}$, we have also two $G^{F}$-conjugacy classes of $F$-stable maximal tori whose representatives can be chosen as $T^{\prime}_{w}:=T_{w}\cap\operatorname{SL}_{n}$ with $w=1,\sigma$.
The following table gives the values of Deligne-Lusztig characters $R_{T_{w}}^{G}(\theta)$ in the case of $G=\operatorname{GL}_{2}$ (and by restriction we get also the values in the $G=\operatorname{SL}_{2}$ case), see [9, §15.9]
$\ &\left(\begin{array}[]{cc}a&0\\ 0&a\end{array}\right)&\left(\begin{array}[]{cc}a&0\\ 0&b\end{array}\right)&\left(\begin{array}[]{cc}x&0\\ 0&x^{q}\end{array}\right)&\left(\begin{array}[]{cc}a&1\\ 0&a\end{array}\right)\\ &a\in\mathbb{F}_{q}^{*}&a\neq b\in\mathbb{F}_{q}^{*}&x\neq x^{q}\in\mathbb{F}_{q^{2}}^{*}&a\in\mathbb{F}_{q}^{*}\\ \hline\cr&&&&\\ R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha,\beta)&(q+1)\alpha(a)\beta(a)&\alpha(a)\beta(b)+\alpha(b)\beta(a)&0&\alpha(a)\beta(a)\\ \alpha,\beta\in\widehat{\mathbb{F}_{q}^{*}}&&&&\\ \hline\cr&&&&\\ R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega)&(1-q)\omega(a)&0&\omega(x)+\omega(x^{q})&\omega(a)\\ \omega\in\widehat{\mathbb{F}_{q^{2}}^{*}}&&&&\\ \hline\cr\end{array}$ (4.1.1)
The character tables of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ and $\operatorname{SL}_{2}(\mathbb{F}_{q})$ from Deligne-Lusztig theory can be found for instance in [9, 15.9].
List of the irreducible characters of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ :
(1) If $\alpha\neq\beta\in\widehat{\mathbb{F}_{q}^{*}}$, then the Deligne-Lusztig characters $R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha,\beta)$ are irreducible (from Lusztig-Jordan decomposition they correspond to conjugacy classes of the second column).
(2) If $\omega\neq\omega^{q}\in\widehat{\mathbb{F}_{q^{2}}^{*}}$, then $-R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega)$ is an irreducible character (it corresponds to conjugacy classes of the third column).
(3) The irreducible characters of the form $\alpha\circ{\rm det}$ and $(\alpha\circ{\rm det})\otimes{\rm St}$ where $\alpha\in\widehat{\mathbb{F}_{q}^{*}}$ and where ${\rm St}$ (called the Steinberg character) is the non-trivial irreducible constituent of the character $R_{T_{1}}^{\operatorname{GL}_{2}}({\rm Id})$. They decomposes into Deligne-Lusztig characters as
$\alpha\circ{\rm det}=\frac{1}{2}\left(R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha\circ{\rm det})+R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\alpha\circ{\rm det})\right)$
$(\alpha\circ{\rm det})\otimes{\rm St}=\frac{1}{2}\left(R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha\circ{\rm det})-R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\alpha\circ{\rm det})\right)$
List of the irreducible characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ :
First of all notice that
$T^{\prime}_{1}{{}^{F}}\simeq\mathbb{F}_{q}^{*},\hskip 28.45274ptT^{\prime}_{\sigma}{{}^{F}}\simeq\mu_{q+1}\subset\mathbb{F}_{q^{2}}^{*}.$
(1) If $\alpha\in\widehat{\mathbb{F}_{q}^{*}}$ satisfies $\alpha^{2}\neq 1$, then $R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha)$ is an irreducible character of $\operatorname{SL}_{2}^{F}$.
(2) If $\omega\in\widehat{\mu_{q+1}}$ satisfies $\omega^{2}\neq 1$, then $-R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega)$ is irreducible.
(3) We have the unipotent characters ${\rm Id}$ and ${\rm St}$ which are the restrictions of the unipotent characters of $\operatorname{GL}_{2}(\mathbb{F}_{q})$.
(4) If $\alpha_{o}$ is the nontrivial square of $1$ in $\widehat{\mathbb{F}_{q}^{*}}$, then $R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})$ splits as a sum of two irreducible characters $\chi_{\alpha_{o}}^{+}$ and $\chi_{\alpha_{o}}^{-}$ of same degree $\frac{q+1}{2}$.
(5) If $\omega_{o}$ is the non-trivial square of $1$ in $\widehat{\mu_{q+1}}$, then $-R_{T_{\sigma}^{\prime}}^{\operatorname{SL}_{2}}(\omega_{o})$ splits as a sum of two irreducible characters $\chi_{\omega_{o}}^{+}$ and $\chi_{\omega_{o}}^{-}$ both of same degree $\frac{q-1}{2}$.
All irreducible characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ in cases (1), (2) and (3) arise as the restriction of irreducible characters of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ and so can be written as a linear combination of Deligne-Lusztig characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$. The irreducible characters $\chi_{\alpha_{o}}^{\pm}$ and $\chi_{\omega_{o}}^{\pm}$ are not restrictions of irreducible characters of $\operatorname{GL}_{2}$ and can not be written as a linear combination of Deligne-Lusztig characters of $\operatorname{SL}_{2}$.
Recall that we have an exact sequence
[…]
where $N:\mathbb{F}_{q^{2}}^{*}\rightarrow\mathbb{F}_{q}^{*},x\mapsto x^{q+1}$ is the norm map.
This dualizes into an exact sequence
[…]
where ${\rm res}$ is the restriction.
The embedding $N^{*}$ identifies $\widehat{\mathbb{F}_{q}^{*}}$ with the linear characters $\omega$ of $\mathbb{F}_{q^{2}}^{*}$ such that $\omega^{q}=\omega$. Therefore, any $\omega\in{\rm res}^{-1}(\omega_{o})$ satisfies $\omega^{q}\neq\omega$.
We thus have the following lemma.
Lemma 4.1.2.
The (non-irreducible) characters $R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})$ and $-R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o})$ are the restrictions respectively of the irreducible characters $R_{T_{1}}^{\operatorname{GL}_{2}}(1,\alpha_{o})$ and $-R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega)$ (for $\omega\in{\rm res}^{-1}(\omega_{o})$).
We will see in the next sections that all the irreducible characters of $\operatorname{GL}_{2}$ and those of $\operatorname{SL}_{2}$ in cases (1), (2) and (3) have a geometric nature which is not the case of the irreducible characters of $\operatorname{SL}_{2}$ in case (4). More precisely Lusztig has defined certain simple perverse sheaves (called character-sheaves) on $\operatorname{SL}_{2}(\overline{\mathbb{F}}_{q})$ whose characteristic functions coincide with the irreducible characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ in cases (1), (2) and (3) but differs from the irreducible characters in case (4) and (5).
4.2 Characteristic functions of orbital complexes
In what follows, $X$ is an algebraic variety over $\overline{\mathbb{F}}_{q}$ equipped with a geometric Frobenius $F:X\rightarrow X$ (or equivalently, an $\mathbb{F}_{q}$-scheme $X_{o}$ such that $X=X_{o}\times_{\mathbb{F}_{q}}\overline{\mathbb{F}}_{q}$), for more details see for instance [9, Chapter 4].
Let $D^{b}_{c}(X)$ be the category of $\overline{{\mathbb{Q}}}_{\ell}$-complexes with constructible cohomology. An $F$-equivariant structure (or Weil structure) on $\mathcal{F}\in D^{b}_{c}(X)$ is an isomorphism
$\varphi:F^{*}(\mathcal{F})\to\mathcal{F}.$
Remark 4.2.1.
If $\mathcal{F}$ is the pullback of a complex $\mathcal{F}_{o}$ on $X_{o}$, then it admits a canonical $F$-equivariant structure, see for instance [14, Chapter 1].
We then say that $(\mathcal{F},\varphi)$ is an $F$-equivariant complex on $X$. Given two $F$-equivariant complexes $(\mathcal{F},\varphi)$ and $(\mathcal{F}^{\prime},\varphi^{\prime})$, the Frobenius $F$ acts on ${\rm Hom}(\mathcal{F},\mathcal{F}^{\prime})$ as
$f\mapsto\varphi^{\prime}\circ F^{*}(f)\circ\varphi^{-1}.$
We denote by $D_{c}^{b}(X;F)$ the category of $F$-equivariant complexes on $X$ with ${\rm Hom}(\mathcal{F},\mathcal{F^{\prime}})^{F}$ as set of morphisms $(\mathcal{F},\varphi)\rightarrow(\mathcal{F}^{\prime},\varphi^{\prime})$.
The characteristic function of $(\mathcal{F},\varphi)\in D_{c}^{b}(X;F)$ is the function $\operatorname{{\bf X}}_{\mathcal{F},\varphi}:X^{F}\to\overline{{\mathbb{Q}}}_{\ell}$ defined as
$\operatorname{{\bf X}}_{\mathcal{F},\varphi}(x)\coloneqq\sum_{i\in\mathbb{Z}}(-1)^{i}\operatorname{Tr}(\varphi_{x}^{i}:\mathcal{H}^{i}_{x}(\mathcal{F})\to\mathcal{H}^{i}_{x}(\mathcal{F})).$
The function $\operatorname{{\bf X}}_{\mathcal{F},\varphi}$ does depend on the choice of the isomorphism $\varphi$. However, in all the cases of relevance for this article, we can make a canonical choice of the isomorphism $\varphi$ and we will drop it from the notation. In particular, if $X$ is an algebraic group, we will always assume that $\varphi_{e}=Id$.
Assume that $G$ is a connected linear algebraic group over $\overline{\mathbb{F}}_{q}$ equipped with a geometric Frobenius $F:G\rightarrow G$. For an $F$-stable conjugacy class $C$ of $G$ together with an $F$-stable $G$-equivariant irreducible local system $\mathcal{E}$ on $C$, we denote by $IC^{\bullet}_{\overline{C},\mathcal{E}}$ the intersection cohomology complex on the Zariski closure $\overline{C}$ of $C$ with coefficients in $\mathcal{E}$. We also fix an $F$-equivariant structure $\varphi:F^{*}(\mathcal{E})\simeq\mathcal{E}$ and we denote again by $\varphi$ the induced $F$-equivariant structure on $IC^{\bullet}_{\overline{C},\mathcal{E}}$ (as $IC^{\bullet}_{\overline{C},\mathcal{E}}$ is irreducible, any two $F$-equivariant structure differ by a scalar).
Proposition 4.2.2.
[17, Proposition 4.4.13] The set $\{{\bf X}_{IC^{\bullet}_{\overline{C},\mathcal{E}},\varphi}\}$, where $(C,\mathcal{E})$ runs over the pairs as above, forms a basis of the space $\mathcal{C}(G^{F})$ of class functions $G^{F}/G^{F}\rightarrow\overline{{\mathbb{Q}}}_{\ell}$.
The above basis is a geometric counterpart of the basis of characteristic functions of conjugacy classes of $G^{F}$.
When $G$ is reductive, the characteristic functions of the $F$-stable character-sheaves will be the geometric counter-part of the basis of irreducible characters of $G^{F}$.
4.3 Character-sheaves on tori
Let $T$ be a torus defined over $\mathbb{F}_{q}$, with geometric Frobenius $F:T\to T$. A Kummer local system $\mathcal{E}$ is a $\overline{{\mathbb{Q}}}_{\ell}$-local system on $T$ such that $\mathcal{E}^{\otimes m}\simeq\overline{{\mathbb{Q}}}_{\ell}$ for some $m\in\mathbb{N}$ such that $(m,q)=1$. Notice that in particular every Kummer local system is of rank $1$ and thus simple.
For any $F$-stable Kummer local system $\mathcal{E}$ on $T$, the characteristic function $\operatorname{{\bf X}}_{\mathcal{E}}$ is a linear character of the finite group $T^{F}$ and any linear character of $T^{F}$ is obtained in this way, i.e.
Proposition 4.3.1.
[25, Proposition 2.3.1] The map $\mathcal{E}\to\operatorname{{\bf X}}_{\mathcal{E}}$ defines a bijection between the set of $F$-stable isomorphism classes of Kummer local systems on $T$ and the set $\widehat{T^{F}}$ of linear characters of $T^{F}$.
The Kummer local systems are the character-sheaves for $T$.
Example 4.3.2.
Consider $T=\mathbb{G}_{m}$ with the canonical Frobenius $F(x)=x^{q}$ for $x\in\mathbb{G}_{m}$. In this case, we have $T^{F}=\mathbb{F}_{q}^{*}$. Consider a linear character $\alpha:\mathbb{F}_{q}^{*}\to\mathbb{C}^{*}$ and let $n$ be the order of $\alpha$. In particular, $n$ divides $q-1$. Fix a surjection $q_{n}:\mathbb{F}_{q}^{*}\to\mathbb{Z}/n\mathbb{Z}$ (by sending a generator $\zeta$ of the cyclic group $\mathbb{F}_{q}^{*}$ to its subgroup of order $n$ generated by $\zeta^{\frac{q-1}{n}}$). Since $\alpha^{n}=1$, there exists a linear character $\mu:\mathbb{Z}/n\mathbb{Z}\to\mathbb{C}^{*}$ such that $\mu\circ q_{n}=\alpha$.
Consider now the $\mathbb{Z}/n\mathbb{Z}$-Galois cover $f_{n}:\mathbb{G}_{m}\to\mathbb{G}_{m}$ given by $f_{n}(z)=z^{n}$. We have a splitting
$(f_{n})_{*}(\overline{{\mathbb{Q}}}_{\ell})=\bigoplus_{\xi\in\widehat{\mathbb{Z}/n\mathbb{Z}}}\mathcal{E}_{\xi}.$
Since $f_{n}$ commutes with $F$, the local systems $\mathcal{E}_{\xi}$ are defined over $\mathbb{F}_{q}$ and have thus a canonical $F$-equivariant structure, and we have $\operatorname{{\bf X}}_{\mathcal{E}_{\mu}}=\alpha$.
4.4 Grothendieck-Springer resolution and geometric induction
For more details and references on this section, see [15, §6.1, 6.2] and also [16, §2.4, 2.7].
Consider a connected reductive algebraic group $G$ over $\overline{\mathbb{F}}_{q}$, a maximal torus $T\subseteq G$ and a Borel subgroup $B\supseteq T$. Let $\pi:B\to T$ be the canonical projection and $W=W_{G}(T)$ be the Weyl group of $G$ with respect to $T$. We denote by $\widetilde{G}$ the variety
$\widetilde{G}\coloneqq\{(xB,g)\in(G/B)\times G\ |\ x^{-1}gx\in B\}.$
We denote by $p:\widetilde{G}\to G$ and $q:\widetilde{G}\to T$ the maps given by $p((xB,g))=g$ and $q((xB,g))=\pi(x^{-1}gx)$. For any $\mathcal{F}\in D^{b}_{c}(T)$, we define the geometric induction $\mathcal{I}_{T}^{G}(\mathcal{F})\in D^{b}_{c}(G)$ as
$\mathcal{I}_{T}^{G}(\mathcal{F})\coloneqq p_{!}q^{*}(\mathcal{F})[{\rm dim}\,G-{\rm dim}\,T].$
By [3], the functor $\mathcal{I}_{T}^{G}$ preserves perverse sheaves and induces thus a functor $\mathcal{I}_{T}^{G}:\mathcal{M}(T)\rightarrow\mathcal{M}(G)$ between categories of perverse sheaves.
Consider the morphism of correspondences
[…]
The morphism $(q,p)$ is small and so $(q,p)_{!}\overline{{\mathbb{Q}}}_{\ell}$ is the intersection cohomology complex $IC^{\bullet}_{S,\overline{{\mathbb{Q}}}_{\ell}}$ with coefficients in the constant sheaf $\overline{{\mathbb{Q}}}_{\ell}$.
From the projection formula we deduce that
$\mathcal{I}_{T}^{G}(\mathcal{F})={\rm pr}_{2\,!}\left({\rm pr}_{1}^{*}(\mathcal{F})\otimes IC^{\bullet}_{S,\overline{{\mathbb{Q}}}_{\ell}}\right).$
Notice also that ${\rm pr}_{1}$ and ${\rm pr}_{2}$ are $W$-equivariant for the natural action of $W$ on $T$ and the trivial action of $W$ on $G$. Therefore, if $\mathcal{F}$ is equipped with a $W$-equivariant structure, namely if we are given a collection $\theta=(\theta_{w})_{w\in W}$ of isomorphisms
$\theta_{w}:w^{*}(\mathcal{F})\rightarrow\mathcal{F}$
such that $\theta_{ww^{\prime}}=\theta_{w}\circ w^{*}(\theta_{w^{\prime}})$ and $\theta_{1}={\rm Id}$, then we get an action of $W$ on $\mathcal{I}_{T}^{G}(\mathcal{F})$, i.e. we have a group homomorphism
$\theta^{G}:W\rightarrow{\rm Aut}(\mathcal{I}_{T}^{G}(\mathcal{F})).$
Remark 4.4.1.
Notice that $W$ does not act on $\widetilde{G}$ so, to construct $\theta^{G}$, we need the factorization through $S$ (Stein factorization) together with the fact that $IC^{\bullet}_{S,\overline{{\mathbb{Q}}}_{\ell}}$ is naturally $W$-equivariant.
Therefore if $\mathcal{F}$ is equipped with a $W$-equivariant structure $\theta$, then we have a decomposition
$\mathcal{I}_{T}^{G}(\mathcal{F})=\bigoplus_{\chi\in\widehat{W}}\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ (4.4.2)
where $\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}\rightarrow\mathcal{I}_{T}^{G}(\mathcal{F})$ is the kernel of the idempotent $1-e_{\chi}\in{\rm End}\left(\mathcal{I}_{T}^{G}(\mathcal{F})\right)$ with
$e_{\chi}$ $=\frac{\chi(1)}{|W|}\sum_{w\in W}\overline{\chi(w)}\,\theta^{G}(w)$
$=\frac{\chi(1)}{|W|}\sum_{w\in W}\chi(w)\,\theta^{G}(w)$
(the last equality follows from the fact that the irreducible characters of Weyl groups have integer values).
Now we assume that $T$ is $F$-stable and that $F$ acts trivially on $W$ (this is will be our case since we will consider $\operatorname{GL}_{2}$, $\operatorname{SL}_{2}$ and $\operatorname{PGL}_{2}$). We do not assume that $B$ is $F$-stable.
The morphisms ${\rm pr}_{1}$ and ${\rm pr}_{2}$ are $F$-equivariant as well as the complex $IC^{\bullet}_{S,\overline{{\mathbb{Q}}}}$. Therefore the functor $\mathcal{I}_{T}^{G}$ induces functors $D_{c}^{b}(T;F)\rightarrow D_{c}^{b}(G;F)$ and $\mathcal{M}(T;F)\rightarrow\mathcal{M}(G;F)$ between $F$-equivariant objects.
Consider $\mathcal{F}\in D_{c}^{b}(T)$ equipped with both a $W$-equivariant structure $\theta$ and an $F$-equivariant structure $\varphi:F^{*}(\mathcal{F})\simeq\mathcal{F}$ such that the two are compatible, namely the following diagram commutes for all $w\in W$, see [15, Diagram (6.1)]:
[…]
Then we have the following compatibility
[…]
For any $\chi\in\widehat{W}$, we see from this compatibility condition (see [15, Remark 6.1.3] for more details) that the canonical $F$-equivariant structure $\varphi^{G}:F^{*}\left(\mathcal{I}_{T}^{G}(\mathcal{F})\right)\simeq\mathcal{I}_{T}^{G}(\mathcal{F})$ restricts to an $F$-equivariant structure $\varphi^{G}_{\chi}:F^{*}(\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi})\simeq\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ and
${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\varphi^{G}}=\sum_{\chi\in\widehat{W}}{\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\varphi^{G}_{\chi}}.$
We have also a formula for ${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\theta^{G}(w)\circ\varphi^{G}}$ for any $w\in W$ which simplifies when the group $W$ is commutative (which will be our case), i.e.
Proposition 4.4.2.
Assume that $G=\operatorname{GL}_{2},\operatorname{SL}_{2}$ or $\operatorname{PGL}_{2}$. Then
${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\theta^{G}(w)\circ\varphi^{G}}=\sum_{\chi\in\widehat{W}}\chi(w)\,{\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\varphi^{G}_{\chi}}$
for all $w\in W$.
Proof.
$\theta^{G}$ induces an action of $W$ on $\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ for all $\chi\in\widehat{W}$ which we also denote by $\theta^{G}$. Then
${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\theta^{G}(w)\circ\varphi^{G}}=\sum_{\chi\in\widehat{W}}{\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\theta^{G}(w)\circ\varphi^{G}_{\chi}}$
We thus need to see that
${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\theta^{G}(w)\circ\varphi^{G}_{\chi}}=\chi(w)\,{\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\varphi^{G}_{\chi}}.$ (4.4.3)
But this follows from the fact that for a representation
$\rho:H\rightarrow\operatorname{GL}(V)$
of a commutative group $H$ in a finite-dimensional complex vector space $V$, then $H$ acts on $V_{\chi}={\rm Ker}(1-e_{\chi})$, with $e_{\chi}=\frac{1}{|H|}\sum_{h\in H}\overline{\chi(h)}\rho(h)$, by the character $\chi$. ∎
Remark 4.4.3.
Note that Formula (4.4.3) for $w=1$ is obviously false if $\chi$ is of degree $>1$. Hence the proof works for commutative $W$ only. For general $W$ we need to decompose further the complexes $\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi}$ which may not be indecomposable.
From the orthogonality relation on irreducible characters of $W$ we deduce.
Proposition 4.4.4.
Under the assumption of Proposition 4.4.2 we have
${\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F})_{\chi},\varphi^{G}_{\chi}}=\frac{1}{|W|}\sum_{w\in W}\chi(w)\,{\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\theta^{G}(w)\circ\varphi^{G}}.$
4.5 Character-sheaves on $\mathrm{SL}_{2}$
Let us start with some generalities: $G$ is an arbitrary connected reductive algebraic group with Frobenius $F$ and $T$ is a maximal torus of $G$.
Theorem 4.5.1 (Lusztig).
Let $\mathcal{L}$ be a Kummer local system (character-sheaf up to a shift) on $T$, then $\mathcal{I}_{T}^{G}(\mathcal{L}[{\rm dim}\,T])$ is a semisimple perverse sheaf on $G$.
Proof.
See for instance [31, §4]. ∎
If $\mathcal{L}$ is a Kummer local system on $T$, then the irreducible constituents of $\mathcal{I}_{T}^{G}(\mathcal{L}[{\rm dim}\,T])$ are examples of character-sheaves on $G$. When $G=\operatorname{GL}_{n}$, we get all character-sheaves in this way. We will see below that this is not true for $\operatorname{SL}_{2}$.
Assume that $T$ is $F$-stable. Since the morphism ${\rm pr}_{2}$ of Diagram (4.4.1) is $G$-equivariant, for any $\mathcal{F}\in D_{c}^{b}(T)$, the complex $\mathcal{I}_{T}^{G}(\mathcal{F})$ is $G$-equivariant and so $\mathcal{I}_{T}^{G}:D_{c}^{b}(T;F)\rightarrow D_{c}^{b}(G;F)$ induces a $\overline{{\mathbb{Q}}}_{\ell}$-linear map
$I_{T}^{G}:\mathcal{C}(T^{F})\rightarrow\mathcal{C}(G^{F}),\hskip 14.22636ptI_{T}^{G}({\bf X}_{\mathcal{F},\varphi})={\bf X}_{\mathcal{I}_{T}^{G}(\mathcal{F}),\varphi^{G}}.$
We have the following particular case of the main result of [23].
Theorem 4.5.2 (Lusztig).
The geometric induction coincides with the Deligne-Lusztig induction, i.e. we have
$I_{T}^{G}=R_{T}^{G}.$
Thanks to the above theorem, the characteristic functions of the $F$-stable character-sheaves are closely related to the irreducible characters of $G^{F}$. When $G=\operatorname{GL}_{n}$, the characteristic functions of the $F$-stable character-sheaves are exactly (up to a sign) the irreducible characters of $\operatorname{GL}_{n}(\mathbb{F}_{q})$. This is not true in general (as we will see in details with $\operatorname{SL}_{2}$). Notice that there is in fact no reason that the decomposition of $\mathcal{I}_{T}^{G}(\mathcal{L})$ as a direct sum of simple objects in the category of perverse sheaves would agree with the decomposition of the character $R_{T}^{G}({\bf X}_{\mathcal{L}})$ as a direct sum of irreducible characters.
We assume now that $G=\operatorname{SL}_{2}$ with $W=S_{2}=\{1,\sigma\}$. The $F$-stable maximal torus $T$ is either the maximal torus of diagonal matrices $T_{1}$ or the twisted $F$-stable maximal torus $T_{\sigma}$ (see §4.1). Recall that $\sigma\in W$ acts by $x\mapsto x^{-1}$ on $T\simeq\operatorname{GL}_{1}$.
Fix a an $F$-stable Kummer local system $\mathcal{L}$.
Remark 4.5.3.
Notice that if $\mathcal{L}$ is $W$-equivariant on $T=T_{1}$, then $\mathcal{L}$ will be also $\sigma F$-stable. Since the pairs $(T_{1},\sigma F)$ is $G$-conjugate to $(T_{\sigma},F)$, we get an $F$-stable Kummer local system $\mathcal{L}_{\sigma}$ on $T_{\sigma}$ and we have
$\mathcal{I}_{T}^{G}(\mathcal{L}[1])\simeq\mathcal{I}_{T_{\sigma}}^{G}(\mathcal{L}_{\sigma}[1])$
as perverse sheaves (not as $F$-equivariant perverse sheaves). To avoid redundancy in the list of character-sheaves we will thus consider only the torus $T_{1}$ in this situation.
We therefore have the following cases :
(1) $T=T_{1}$ or $T_{\sigma}$ and $\mathcal{L}^{\otimes 2}\not\simeq\overline{{\mathbb{Q}}}_{\ell}$, then $\sigma^{*}(\mathcal{L})\not\simeq\mathcal{L}$. The perverse sheaf $\mathcal{I}_{T}^{G}(\mathcal{L}[{\rm dim}\,T])$ is irreducible and so is a character-sheaf.
(2) $T=T_{1}$ and $\mathcal{L}=\overline{{\mathbb{Q}}}_{\ell}$, then $\mathcal{L}$ is naturally $W$-equivariant and the decomposition (4.4.2) reads
$\mathcal{I}_{T}^{G}(\overline{{\mathbb{Q}}}_{\ell}[1])\simeq\overline{{\mathbb{Q}}}_{\ell}[3]\oplus{\rm St}[3]$
where ${\rm St}[3]$ is the character-sheaf corresponding to the sign character of $W$.
(3) $T=T_{1}$ and $\mathcal{L}^{\otimes 2}\simeq\overline{{\mathbb{Q}}}_{\ell}$ with $\mathcal{L}\not\simeq\overline{{\mathbb{Q}}}_{\ell}$, then, by Example 4.3.2, $\mathcal{L}$ is the non-trivial constituent $\mathcal{A}_{o}$ of
$(f_{2})_{*}(\overline{{\mathbb{Q}}}_{\ell})=\overline{{\mathbb{Q}}}_{\ell}\oplus\mathcal{A}_{o}.$
It corresponds to the character $\alpha_{o}$ of $\mathbb{F}_{q}^{*}\simeq T_{1}^{F}$ (which take the value $1$ at squares and $-1$ at non-squares). Since $f_{2}$ is $W$-equivariant, the local system $\mathcal{A}_{o}$ must be $W$-equivariant and by (4.4.2) we have
$\mathcal{I}_{T}^{G}(\mathcal{A}_{o}[1])=\mathcal{X}_{1}[3]\oplus\mathcal{X}_{\epsilon}[3]$
where $X_{\epsilon}$ is the constituent corresponding to the sign character of $W$.
Notice that we do not get all character-sheaves here, we are missing the two cuspidals which are the two non-trivial $G$-equivariant irreducible local systems on the conjugacy classes of
$\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right),\hskip 28.45274pt\text{and}\hskip 28.45274pt\left(\begin{array}[]{cc}-1&1\\ 0&-1\end{array}\right)$
In the first two cases (1) and (2), the characteristic functions of the character-sheaves coincide with the irreducible "characters" $R_{T}^{G}({\bf X}_{\mathcal{L}})$, ${\rm Id}$ and ${\rm St}$ of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ (here by "character" we mean up to a sign, since $R_{T}^{G}(\theta)$ is the negative of a character when $T$ is the twisted torus). This is not true in the last case as we now see.
Denote by $X_{1}$ and $X_{\epsilon}$ the characteristic functions of $\mathcal{X}_{1}$ and $\mathcal{X}_{\epsilon}$.
Theorem 4.5.4.
$X_{1}$ $=\frac{1}{2}\left(R_{T_{1}}^{G}(\alpha_{o})+R_{T_{\sigma}}^{G}(\omega_{o})\right)$
$X_{\epsilon}$ $=\frac{1}{2}\left(R_{T_{1}}^{G}(\alpha_{o})-R_{T_{\sigma}}^{G}(\omega_{o})\right)$
Proof.
Proposition 4.4.4 says that
$X_{1}$ $=\frac{1}{2}\left({\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\varphi^{G}}+{\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\theta^{G}(\sigma)\circ\varphi^{G}}\right)$
$X_{\epsilon}$ $=\frac{1}{2}\left({\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\varphi^{G}}-{\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\theta^{G}(\sigma)\circ\varphi^{G}}\right)$
By Theorem 4.5.2 applied to $T=T_{1}$ we have
${\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\varphi^{G}}=R_{T_{1}}^{G}(\alpha_{o}).$
As in Remark 4.5.3, the $F$-equivariant local system $\mathcal{A}_{o}$ (affording the character $\alpha_{o}$ on $T_{1}^{F}$) being naturally $W$-equivariant (with $W$-equivariant structure compatible with $F$-equivariant structure) is transported to an $F$-equivariant local system $\mathcal{A}_{\sigma}$ on $T_{\sigma}$ which corresponds to the non-trivial square $\omega_{o}$ in $\widehat{\mu_{q+1}}$ of the identity character.
Therefore by Theorem 4.5.2 applied to $T=T_{\sigma}$ we have
${\bf X}_{\mathcal{I}_{T_{1}}^{G}(\mathcal{A}_{o}),\theta^{G}(\sigma)\circ\varphi^{G}}$ $={\bf X}_{\mathcal{I}_{T_{\sigma}}^{G}(\mathcal{A}_{\sigma})}$
$=R_{T_{\sigma}}^{G}(\omega_{o})$
(the first equality being formal). ∎
Now notice that
Lemma 4.5.5.
$\omega_{o}(-1)=-\alpha_{o}(-1).$
Proof.
This follows from the fact that $-1$ is a square in $\mathbb{F}_{q}^{*}$ if and only if it is a non-square in $\mu_{q+1}\subset\mathbb{F}_{q^{2}}^{*}$. ∎
From Theorem 4.5.4, Lemma 4.1.2 and the Table (4.1.1) we deduce the following table for the values of $X_{1}$ and $X_{\epsilon}$:
$\ &\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)&\left(\begin{array}[]{cc}-1&0\\ 0&-1\end{array}\right)&\left(\begin{array}[]{cc}a&0\\ 0&a^{-1}\end{array}\right)&\left(\begin{array}[]{cc}x&0\\ 0&x^{q}\end{array}\right)&\left(\begin{array}[]{cc}1&1\\ 0&1\end{array}\right)&\left(\begin{array}[]{cc}-1&1\\ 0&-1\end{array}\right)\\ &&&a\neq a^{-1}\in\mathbb{F}_{q}^{\times}&x^{q+1}=1,x\neq x^{q}&&\\ \hline\cr&&&&&&\\ X_{1}&1&q\alpha_{o}(-1)&\alpha_{o}(a)&0&1&0\\ &&&&&&\\ \hline\cr&&&&&&\\ X_{\epsilon}&q&\alpha_{o}(-1)&0&\omega_{o}(x)&1&\alpha_{o}(-1)\\ &&&&&&\\ \hline\cr\end{array}$ (4.5.1)
We saw in §4.1, that the irreducible characters $\chi_{\alpha_{o}}^{\pm}$ and $\chi_{\omega_{o}}^{\pm}$ have degrees $\frac{q+1}{2}$ and $\frac{q-1}{2}$ and so $X_{1}$ and $X_{\epsilon}$ can not be irreducible characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$.
5 Main results
5.1 Statements
Let $K$ be an algebraically closed field of characteristic $\neq 2$ ($K=\mathbb{C}$ or $\overline{\mathbb{F}}_{q}$) and $\mathcal{C}=(\mathcal{C}_{1},\dots,\mathcal{C}_{k})$ be a $k$-tuple of regular semisimple conjugacy classes of $\operatorname{PGL}_{2}(K)$. When $K=\overline{\mathbb{F}}_{q}$, we assume that $\mathcal{C}$ respects the conditions explained at the beginning of section 3.4. We denote by $\mathcal{M}_{C}(K)$ the corresponding character stack.
The non-degenerate case
We assume that all conjugacy classes $\mathcal{C}_{i}$ are non-degenerate (i.e. that their stabilisers are connected).
We fix a generic $k$-tuple $R^{\prime}=(R^{\prime}_{1},\dots,R^{\prime}_{k})$ of irreducible characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ of the form
$R^{\prime}=\left(R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{1}),\dots,R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{k})\right)$
(see §5.4 for the definition of generic $R^{\prime}$).
Theorem 5.1.1.
We have
$E(\mathcal{M}_{\mathcal{C}}(\mathbb{C});q)=E(\mathcal{M}_{\mathcal{C}}(\overline{\mathbb{F}}_{q});q)=\left\langle 1_{(\mathcal{C}_{1})^{F}}*\cdots*1_{(\mathcal{C}_{k})^{F}},1_{\{1\}}\right\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}.$ (5.1.1)
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k},1\right\rangle_{\operatorname{SL}_{2}(\mathbb{F}_{q})}=2\,\mathbb{A}_{0}(0,\sqrt{q}).$ (5.1.2)
where $\mathbb{A}_{r}(z,w)$ is defined in §3.4.
We will see that these polynomials do not depend on the choice of the eigenvalues as long as this choice is generic and the conjugacy classes regular non-degenerate.
The proof of the first assertion is completely similar to the analogous statement for $\operatorname{GL}_{n}(\mathbb{C})$-character varieties (see [11]). Namely we first notice that the RHS counts the number of points of $\mathcal{M}_{\mathcal{C}}$ over $\mathbb{F}_{q}$ and we conclude by Theorem 2.1.1.
The second assertion will be a particular case of Theorem 5.4.4(2) with $k_{2}=0$.
It follows from Conjecture 3.4.3 and Remark 3.4.4 that
${H_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C});q,t)}$${\langle 1_{(\mathcal{C}_{1})^{F}}*\cdots*1_{(\mathcal{C}_{k})^{F}},1_{1}\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}}$${q^{k-3}\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k},1\rangle_{\operatorname{SL}_{2}(\mathbb{F}_{q})}}$$\scriptstyle{t\mapsto-1}$ pure part
The degenerate case
We fix a non-negative integer $1\leq m<k$ and we assume that $\mathcal{C}_{1}=\dots=\mathcal{C}_{m}=\mathcal{C}_{-1}$ and that the $\mathcal{C}_{i}$’s are non-degenerate for $i>m$.
Recall that $\mathcal{L}_{\epsilon}$ denotes the non-trivial irreducible $\operatorname{PGL}_{2}(K)$-equivariant local system on $\mathcal{C}_{-1}$.
If $K=\overline{\mathbb{F}}_{q}$, it is equipped with a natural $F$-equivariant structure and we denote by
$Y_{\epsilon}:\operatorname{PGL}_{2}(\mathbb{F}_{q})\rightarrow\overline{{\mathbb{Q}}}_{\ell}$
the characteristic function of $\mathcal{L}_{\epsilon}$. It is supported on $(\mathcal{C}_{-1})^{F}$.
Remark 5.1.2.
The set $(\mathcal{C}_{-1})^{F}$ is not a single conjugacy class of $\operatorname{PGL}_{2}(\mathbb{F}_{q})$. Since the stabiliser of an element of $\mathcal{C}_{-1}$ has two connected components, it decomposes as
$(\mathcal{C}_{-1})^{F}=\mathcal{O}_{e}\sqcup\mathcal{O}_{\sigma},$
the disjoint union of two $\operatorname{PGL}_{2}(\mathbb{F}_{q})$-conjugacy classes. Here $\mathcal{O}_{e}$ is the $\operatorname{PGL}_{2}(\mathbb{F}_{q})$-conjugacy class of the image $D_{e}=p(g_{-1})$ of the matrix $g_{-1}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$ in $\operatorname{PGL}_{2}$. For the other one we fix $x\in\mathbb{F}_{q^{2}}^{*}$ such that $x^{q}=-x$. The matrix
$\left(\begin{array}[]{cc}x&0\\ 0&-x\end{array}\right)$
is $\sigma F$-stable and so is $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$-conjugate to an $F$-stable matrix $g_{\sigma}$ in $T_{\sigma}$. The image $D_{\sigma}=p(g_{\sigma})$ in $\operatorname{PGL}_{2}$ of this latter matrix is $F$-stable and is geometrically conjugate to $D_{e}$ (but not rationally as $x\notin\mathbb{F}_{q}$). The second conjugacy class $\mathcal{O}_{\sigma}$ is the $\operatorname{PGL}_{2}(\mathbb{F}_{q})$-conjugacy class of $D_{\sigma}$.
Then
$Y_{\epsilon}=1_{\mathcal{O}_{e}}-1_{\mathcal{O}_{\sigma}},$
i.e. it takes the value $1$ at $\mathcal{O}_{e}$ and $-1$ at $\mathcal{O}_{\sigma}$.
Fix a subset $B\subset\{2,\dots,m\}$ (possibly empty) and let $({\bf Y}_{1},\dots,{\bf Y}_{k})$ be the $k$-tuple of functions ${\bf Y}_{i}:\operatorname{PGL}_{2}(\mathbb{F}_{q})\rightarrow\overline{{\mathbb{Q}}}_{\ell}$ given by
${\bf Y}_{i}=\begin{cases}Y_{\epsilon}&\text{ if }i\in\{1\}\cup B,\\ 1_{(\mathcal{C}_{i})^{F}}&\text{ otherwise.}\end{cases}$
where $1_{(\mathcal{C}_{i})^{F}}$ is the characteristic function of $(\mathcal{C}_{i})^{F}$ that takes the value $1$ on $(\mathcal{C}_{i})^{F}$ and 00 elsewhere.
Theorem 5.1.3.
We have
$E(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{F}_{B};q)=E(\mathcal{M}_{\mathcal{C}}(\overline{\mathbb{F}}_{q}),\mathcal{F}_{B};q)=\left\langle{\bf Y}_{1}*\cdots*{\bf Y}_{k},1_{\{1\}}\right\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}$
Consider a generic $k$-tuple $({\mathbb{X}}_{1},\dots,{\bf X}_{k})$ of characteristic functions of character-sheaves on $\operatorname{SL}_{2}$ (for the definition of genericity, see in §5.5) such that
${\bf X}_{i}=\begin{cases}X_{\epsilon}&\text{ if }i\in\{1\}\cup B,\\ X_{1}&\text{ if }i\in\{2,\dots,m\}\backslash B,\\ R_{T^{\prime}_{2}}^{\operatorname{SL}_{2}}(\gamma_{i})&\text{ otherwise.}\end{cases}$
Theorem 5.1.4.
We have
$\left\langle{\bf X}_{1}\cdots{\bf X}_{k},1\right\rangle_{\operatorname{SL}_{2}}=\frac{1}{2^{m-1}}\sum_{\begin{subarray}{c}r=0\\ r\text{ even}\end{subarray}}^{m}C_{|B|+1,m-|B|-1,r}\,\,\mathbb{A}_{r}\left(0,\sqrt{q}\right).$ (5.1.3)
We deduce from the above theorems together with the conjectural formula (3.4.7) the following conjectural picture :
${H_{c}(\mathcal{M}_{\mathcal{C}}(\mathbb{C}),\mathcal{F}_{B};q,t)}$${\langle{\bf Y}_{1}*\cdots*{\bf Y}_{k},1_{1}\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}}$${q^{k-3}\langle{\bf X}_{1}\cdots{\bf X}_{k},1\rangle_{\operatorname{SL}_{2}(\mathbb{F}_{q})}}$$\scriptstyle{t\mapsto-1}$ pure part
We will start with the proof of Theorem 5.1.3.
Notice that the computation of the LHS of (5.1.3) will reduce to some computations in $\operatorname{GL}_{2}$ thanks to the Frobenius reciprocity formula
$\left\langle{\rm Res}^{\operatorname{GL}_{2}(\mathbb{F}_{q})}_{\operatorname{SL}_{2}(\mathbb{F}_{q})}(f),1\right\rangle_{\operatorname{SL}_{2}}=\left\langle f,{\rm Ind}_{\operatorname{SL}_{2}(\mathbb{F}_{q})}^{\operatorname{GL}_{2}(\mathbb{F}_{q})}(1)\right\rangle_{\operatorname{GL}_{2}}$
for any class function $f$ on $\operatorname{GL}_{2}(\mathbb{F}_{q})$.
5.2 Some result for $\mathrm{GL}_{2}$
Consider a $k$-tuple $R=(R_{1},\dots,R_{k})$ of Deligne-Lusztig characters of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ (possible not irreducible) of the form
$R=\left(R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha_{1}^{1},\alpha^{1}_{2}),\dots,R_{T_{1}}^{\operatorname{GL}_{2}}(\alpha^{k_{1}}_{1},\alpha^{k_{1}}_{2}),-R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega_{1}),\dots,-R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega_{k_{2}})\right)$
with $k_{1}>0$ (but we allow $k_{2}$ to be zero).
We have the following definition [19, Definition 6.8.6].
Definition 5.2.1.
We say that $R$ is generic if
$\prod_{i=1}^{k_{1}}(\alpha^{i}_{1}\alpha^{i}_{2})\prod_{j=1}^{k_{2}}(\omega_{j}|_{\mathbb{F}_{q}^{*}})=1.$
(2) If $k_{2}=0$, we further ask that for any $w_{1},\dots,w_{k}\in S_{2}$, we have
$\prod_{i=1}^{k}\alpha^{i}_{w_{i}(1)}\neq 1.$
Theorem 5.2.2.
[19, Theorem 6.10.1] If $R$ is generic then
$\left\langle R_{1}\otimes\cdots\otimes R_{k},1\right\rangle_{\operatorname{GL}_{2}}=(-1)^{k_{2}}\mathbb{A}_{k_{2}}(0,\sqrt{q}).$
Remark 5.2.3.
It follows from the theorem that
$\left\langle R_{1}\otimes\cdots\otimes R_{k},1\right\rangle_{\operatorname{GL}_{2}}$
does not depend on the choice of the characters $(\alpha_{1}^{i},\alpha_{2}^{i})$ and $\omega^{j}$ as long as the choice is generic.
Remark 5.2.4.
If the condition (1) in Definition 5.2.1 is not satisfied, then
$\left\langle R_{1}\otimes\cdots\otimes R_{k},1\right\rangle_{\operatorname{GL}_{2}}=0.$
This is analoguous to the condition (2.2.1) for the character stack $\mathcal{M}_{\mathcal{C}}$ to be non-empty.
5.3 From $\mathrm{GL}_{2}$ to $\mathrm{SL}_{2}$
Consider a $k$-tuple $R^{\prime}=(R^{\prime}_{1},\dots,R^{\prime}_{k})$ of Deligne-Lusztig characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ (possibly not irreducible) of the form
$R^{\prime}=\left(R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{1}),\dots,R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{k_{1}}),-R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\eta_{1}),\dots,-R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\eta_{k_{2}})\right).$
Each $R^{\prime}_{i}$ is the restriction of a Deligne-Lusztig character $R_{i}$ of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ of the form
$R_{i}=\begin{cases}R_{T^{\prime}_{1}}^{\operatorname{GL}_{2}}(1,\gamma_{i})&\text{ if }1\leq i\leq k_{1},\\ -R_{T_{\sigma}}^{\operatorname{GL}_{2}}(\omega_{i})&\text{ if }k_{1}+1\leq i\leq k_{2}.\end{cases}$
By the Frobenius reciprocity formula we have
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k}\right\rangle_{\operatorname{SL}_{2}}=\left\langle R_{1}\otimes\cdots\otimes R_{k},{\rm Ind}_{\operatorname{SL}_{2}}^{\operatorname{GL}_{2}}(1)\right\rangle_{\operatorname{GL}_{2}}.$
Since
${\rm Ind}_{\operatorname{SL}_{2}}^{\operatorname{GL}_{2}}(1)=\sum_{\delta\in\widehat{\mathbb{F}_{q}^{*}}}\delta\circ\det$
we have
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k}\right\rangle_{\operatorname{SL}_{2}}=\sum_{\delta\in\widehat{\mathbb{F}_{q}^{*}}}\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\delta^{-1}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}$
By Remark 5.2.4, the multiplicity
$\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\delta^{-1}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}$
vanishes unless
$\gamma_{1}\cdots\gamma_{k_{1}}(\omega_{1}|_{\mathbb{F}_{q}^{*}})\cdots(\omega_{k_{2}}|_{\mathbb{F}_{q}^{*}})=\delta^{2}.$
We need the following lemma.
Lemma 5.3.1.
Given $\gamma\in\widehat{\mathbb{F}_{q}^{*}}$, the equation
$\delta^{2}=\gamma$
has a solution if and only if $\gamma(-1)=1$. In this case the two solutions differ by the character $\alpha_{o}$.
Proof.
The group homomorphism
$\varphi:\widehat{\mathbb{F}_{q}^{*}}\rightarrow\{-1,1\},\hskip 28.45274pt\gamma\mapsto\gamma(-1)$
is surjective and we see that the square elements of $\widehat{\mathbb{F}_{q}^{*}}$ live in the kernel of $\varphi$. The subgroup ${\rm Ker}(\varphi)$ is thus of index $2$ in $\widehat{\mathbb{F}_{q}^{*}}$ and it contains the subgroup of square elements of $\widehat{\mathbb{F}_{q}^{*}}$ which is also of index $2$ hence the result. ∎
We thus have the following result.
Proposition 5.3.2.
(1) If $\left(\gamma_{1}\cdots\gamma_{k_{1}}(\omega_{1}|_{\mathbb{F}_{q}^{*}})\cdots(\omega_{k_{2}}|_{\mathbb{F}_{q}^{*}})\right)(-1)=-1$, then
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k},1\right\rangle_{\operatorname{SL}_{2}}=0.$
(2) If $\left(\gamma_{1}\cdots\gamma_{k_{1}}(\omega_{1}|_{\mathbb{F}_{q}^{*}})\cdots(\omega_{k_{2}}|_{\mathbb{F}_{q}^{*}})\right)(-1)=1$, then we choose a square root
$\lambda_{R}=\sqrt{\gamma_{1}\cdots\gamma_{k_{1}}(\omega_{1}|_{\mathbb{F}_{q}^{*}})\cdots(\omega_{k_{2}}|_{\mathbb{F}_{q}^{*}})}$
and we have
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k},1\right\rangle_{\operatorname{SL}_{2}}$
$=\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\lambda_{R}^{-1}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}+\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\lambda_{R}^{-1}\alpha_{o}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}.$
5.4 Generic $k$-tuples of semisimple characters of $\mathrm{SL}_{2}(\mathbb{F}_{q})$
Let $R^{\prime}=(R^{\prime}_{1},\dots,R^{\prime}_{k})$ be as in the previous section.
Definition 5.4.1.
The $k$-tuple is said to be generic if we have :
$\prod_{i=1}^{k_{1}}\gamma_{i}(-1)\prod_{j=1}^{k_{2}}\eta_{i}(-1)=1.$
(2) If $k_{2}=0$, we further ask that for all $w_{1},\dots,w_{k}\in W=S_{2}=\{1,\sigma\}$,
${{}^{w_{1}}}\gamma_{1}\cdots{{}^{w_{k}}}\gamma_{k}\neq 1$
where ${{}^{\sigma}}\gamma$ is the character $\gamma^{-1}$.
Proposition 5.4.2.
The $k$-tuple $R^{\prime}$ is generic if and only if the two $(k+1)$-tuples of irreducible characters of $\operatorname{GL}_{2}(\mathbb{F}_{q})$
$R^{+}=(R_{1},\dots,R_{k},\lambda_{R}^{-1}\circ\det),\hskip 14.22636pt\text{ and }\hskip 14.22636ptR^{-}=(R_{1},\dots,R_{k},\lambda_{R}^{-1}\alpha_{o}\circ\det)$
are both generic.
Remark 5.4.3.
If $T$ is an $F$-stable maximal torus of $\operatorname{GL}_{2}$, $\theta$ a character of $T^{F}$ and $\lambda$ a linear character of $\mathbb{F}_{q}^{*}$, then
$R_{T}^{\operatorname{GL}_{2}}(\theta)\otimes(\lambda\circ\det)=R_{T}^{Gl_{2}}(\theta\otimes(\lambda\circ\det)).$
The $(k+1)$-tuple $R^{+}$ is generic if and only if the $k$-tuple $(R^{\lambda}_{1},R_{2},\dots,R_{k})$, where
$R^{\lambda}_{1}=R_{T_{1}}^{\operatorname{GL}_{2}}\left((1,\gamma_{1})\otimes(\lambda^{-1}_{R}\circ\det)\right),$
is generic.
Proof.
From our choice of $\lambda_{R}$, we have that
$\lambda_{R}^{-2}\prod_{i=1}^{k_{1}}\gamma_{i}\prod_{j=1}^{k_{2}}(\omega_{j}|_{\mathbb{F}_{q}^{*}})=(\lambda_{R}\alpha_{o})^{-2}\prod_{i=1}^{k_{1}}\gamma_{i}\prod_{j=1}^{k_{2}}(\omega_{j}|_{\mathbb{F}_{q}^{*}})=1.$
We see thus that the condition $(1)$ of Definition 5.2.1 is verified for the $(k+1)$-tuples $R^{+},R^{-}$. Assume now that $k_{2}=0$ and consider $w_{1},\dots,w_{k}\in W$. Let $I\subseteq\{1,\dots,k\}$ be the subset
$I\coloneqq\{i\in\{1,\dots,k\}\ |\ w_{i}=\sigma\}.$
Condition $(2)$ of Definition 5.2.1 for the $(k+1)$-tuples $R^{+},R^{-}$ and the elements $w_{1},\dots,w_{k}$ is thus
$\lambda_{R}^{-1}\prod_{i\in I}\gamma_{i}\neq 1\ \ \text{ and }\ \ \lambda_{R}^{-1}\alpha_{o}\prod_{i\in I}\gamma_{i}\neq 1.$ (5.4.1)
Notice that eq. (5.4.1) above holds if and only if
$\lambda_{R}^{-2}\prod_{i\in I}\gamma^{2}_{i}\neq 1.$ (5.4.2)
We have now
$\lambda_{R}^{-2}\prod_{i\in I}\gamma^{2}_{i}=\prod_{i\in I}\gamma_{i}\prod_{j\in I^{c}}\gamma_{j}^{-1}=({{}^{w_{1}}}\gamma_{1}\cdots{{}^{w_{k}}}\gamma_{k})^{-1},$
i.e. eq(5.4.1) is equivalent to the Condition $(2)$ of Definition 5.4.1 for the $k$-tuple $R^{{}^{\prime}}$ and the elements $w_{1},\dots,w_{k}$.
Theorem 5.4.4.
If $R^{\prime}$ is generic then
$\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\lambda_{R}^{-1}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}=\left\langle R_{1}\otimes\cdots\otimes R_{k}\otimes(\lambda_{R}^{-1}\alpha_{o}\circ\det),1\right\rangle_{\operatorname{GL}_{2}}$
$\left\langle R^{\prime}_{1}\otimes\cdots\otimes R^{\prime}_{k},1\right\rangle_{\operatorname{SL}_{2}}=2(-1)^{k_{2}}\mathbb{A}_{k_{2}}(0,\sqrt{q}).$
Proof.
The first assertion is a consequence of Proposition 5.4.2, Remark 5.4.3 and Remark 5.2.3. The second assertion is a consequence of the assertion (1) together with Proposition 5.3.2 and Theorem 5.2.2. ∎
5.5 Proof of Theorem 5.1.3
We use the notation of §5.1.
We assume that the $k$-tuple $({\bf X}_{1},\dots,{\bf X}_{k})$ is generic, meaning that the $k$-tuple
$\left(\underbrace{R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o}),\dots,R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})}_{m},{\bf X}_{m+1},\dots,{\bf X}_{k}\right)$
of Deligne-Lusztig characters of $\operatorname{SL}_{2}(\mathbb{F}_{q})$ is generic.
Put $m_{1}=|B|+1$ and $m_{2}=m-m_{1}$. We have
$\left\langle{\bf X}_{1}\otimes\cdots\otimes{\bf X}_{k},1\right\rangle_{\operatorname{SL}_{2}}=\left\langle(X_{\epsilon})^{\otimes m_{1}}\otimes(X_{1})^{\otimes m_{2}}\otimes\bigotimes_{j=m+1}^{k}R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{j}),1\right\rangle_{\operatorname{SL}_{2}}$
From Theorem 4.5.4, we have
$\left\langle{\bf X}_{1}\otimes\cdots\otimes{\bf X}_{k},1\right\rangle_{\operatorname{SL}_{2}}$
$=\frac{1}{2^{m}}\left\langle(R_{T_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})-R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o}))^{\otimes m_{1}}\otimes(R_{T_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})+R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o}))^{\otimes m_{2}}\otimes\bigotimes_{j=m+1}^{k}R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{j}),1\right\rangle_{\operatorname{SL}_{2}}$
$=\frac{1}{2^{m}}\sum_{r=0}^{m}C_{|B|+1,m-|B|-1,r}\left\langle R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o})^{\otimes r}\otimes R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})^{\otimes m-r}\otimes\bigotimes_{j=m+1}^{k}R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{j}),1\right\rangle_{\operatorname{SL}_{2}}$
By the genericity assumption we have
$\alpha_{o}(-1)^{m}(\gamma_{m+1}\cdots\gamma_{k})(-1)=1.$
Since $\omega_{o}(-1)=-\alpha_{o}(-1)$ by Lemma 4.5.5 we deduce that
$\omega_{o}(-1)^{r}\alpha_{o}(-1)^{m-r}(\gamma_{m+1}\cdots\gamma_{k})(-1)=(-1)^{r}.$ (5.5.1)
By Proposition 5.3.2(1) we see that
$\left\langle R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o})^{\otimes r}\otimes R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})^{\otimes m-r}\otimes\bigotimes_{j=m+1}^{k}R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{j}),1\right\rangle_{\operatorname{SL}_{2}}$
vanishes unless $r$ is even.
Therefore
$\left\langle{\bf X}_{1}\otimes\cdots\otimes{\bf X}_{k},1\right\rangle_{\operatorname{SL}_{2}}$
$=\frac{1}{2^{m}}\sum_{\begin{subarray}{c}r=0\\ r\text{ even}\end{subarray}}^{m}C_{|B|+1,m-|B|-1,r}\left\langle R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o})^{\otimes r}\otimes R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})^{\otimes m-r}\otimes\bigotimes_{j=m+1}^{k}R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{j}),1\right\rangle_{\operatorname{SL}_{2}}$
Now notice that the $k$-tuple
$\left(\underbrace{R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o}),\dots,R_{T^{\prime}_{\sigma}}^{\operatorname{SL}_{2}}(\omega_{o})}_{r},\underbrace{R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o}),\dots,R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\alpha_{o})}_{m-r},R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{m+1}),\dots,R_{T^{\prime}_{1}}^{\operatorname{SL}_{2}}(\gamma_{k})\right)$
is generic for all even integer $r$ ($0\leq r\leq m$).
Indeed, if $r=0$, then this is by assumption, if $0<r$ the condition (1) of Definition 5.4.1 is satisfied by Formula (5.5.1) since $r$ is even (the condition (2) also since $k_{2}=r>0$).
We deduce Theorem 5.1.3 from Theorem 5.4.4(2).
5.6 The Euler specialization
In this section, we prove Theorem 5.1.3. We give two proofs : a formal one and a computational one. The second proof is "dual" to the proof of Theorem 5.1.3 given in section 5.5.
5.6.1 A formal proof
Recall that $\mathcal{C}$ is a generic $k$-tuple of conjugacy classes of $\operatorname{PGL}_{2}(\overline{\mathbb{F}}_{q})$ such that $\mathcal{C}_{1}=\cdots=\mathcal{C}_{m}$ are degenerate and $\mathcal{C}_{j}$ is non-degenerate for $j>m$. Fix $C=(C_{1},\dots,C_{k})$ a corresponding $k$-tuple of $F$-stable conjugacy classes of $\operatorname{GL}_{2}(\overline{\mathbb{F}}_{q})$ with the assumptions made at the beginning of section 3.4.
Notice that, for any $(x_{1},\dots,x_{k})\in X_{\mathcal{C}}(\mathbb{F}_{q})$, we have that
${\bf X_{\mathcal{L}_{B}}}(x_{1},\dots,x_{k})={\bf Y}_{1}(x_{1})\cdots{\bf Y}_{k}(x_{k}).$ (5.6.1)
We thus have
$\left\langle{\bf Y}_{1}\ast\cdots\ast{\bf Y}_{k},1_{\{1\}}\right\rangle=\dfrac{1}{|\operatorname{PGL}_{2}(\mathbb{F}_{q})|}\sum_{(x_{1},\dots,x_{k})\in X_{\mathcal{C}}(\mathbb{F}_{q})}{\bf Y}_{1}(x_{1})\cdots{\bf Y}_{k}(x_{k})=$ (5.6.2)
$=\dfrac{1}{|\operatorname{PGL}_{2}(\mathbb{F}_{q})|}\sum_{(x_{1},\dots,x_{k})\in X_{\mathcal{C}}(\mathbb{F}_{q})}{\bf X_{\mathcal{L}_{B}}}(x_{1},\dots,x_{k})=\sum_{z\in\mathcal{M}_{\mathcal{C}(\mathbb{F}_{q})}}{\bf X}_{\mathcal{F}_{B}}(z).$ (5.6.3)
From the last equality of formula (5.6.3) and Grothendieck’s trace formula, we deduce that
$\left\langle{\bf Y}_{1}\ast\cdots\ast{\bf Y}_{k},1_{\{1\}}\right\rangle=\sum_{k}(-1)^{k}\operatorname{Tr}\left(F\,|\,H^{k}_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B})\right)$
Notice that, since the action of $H_{m}$ and $F$ on $H^{*}_{c}(\mathcal{M}_{\mathcal{C}})$ commute, the decompositon (3.4.3) and the inversion formula in the character ring of $H_{m}$ imply that
$\operatorname{Tr}\left(F\,|\,H^{k}_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B})\right)=\frac{1}{|H_{m}|}\sum_{y\in H_{m}}\operatorname{Tr}\left(yF\,|\,H^{k}_{c}(\mathcal{M}^{+}_{C})\right)\,\overline{\chi_{B}}(y)$
for any $k$ and thus
$\sum_{k}(-1)^{k}\operatorname{Tr}\left(F\,|\,H^{k}_{c}(\mathcal{M}_{\mathcal{C}},\mathcal{F}_{B})\right)=\dfrac{1}{2^{m-1}}\sum_{y\in H_{m}}\left(\sum_{k}(-1)^{k}\operatorname{Tr}\left(yF\,|\,H^{k}_{c}(\mathcal{M}_{C}^{+})\right)\right)\overline{\chi_{B}}(y)$ (5.6.4)
Theorem 5.1.3 is now a consequence of Theorem 2.3.3 together with Grothendieck’s trace formula and Formula (3.4.5).
5.6.2 A computational proof
We give here another proof of Theorem 5.1.3, by direct computations of convolution products of class functions of $\operatorname{PGL}_{2}(\mathbb{F}_{q})$.
We start by the following more general remark. Consider $k$ semisimple conjugacy classes of $\operatorname{PGL}_{2}(\mathbb{F}_{q})$ denoted by $\mathcal{O}_{1},\dots,\mathcal{O}_{k}$ and corresponding conjugacy classes $O_{1},\dots,O_{k}$ of $\operatorname{GL}_{2}(\mathbb{F}_{q})$ such that $\pi(O_{j})=\mathcal{O}_{j}$. We assume that $\mathcal{O}_{1},\dots,\mathcal{O}_{m}$ are degenerate (i.e. $\mathcal{O}_{i}=\mathcal{O}_{e}$ or $\mathcal{O}_{i}=\mathcal{O}_{\sigma}$ for $i=1,\dots,m$) and $\mathcal{O}_{m+1},\dots,\mathcal{O}_{k}$ are non-degenerate. If $\mathcal{O}_{j}=\mathcal{O}_{e}$, we take as $O_{j}$ the conjugacy class of $g_{-1}$ and if $\mathcal{O}_{j}=\mathcal{O}_{\sigma}$ we take as $O_{j}$ the conjugacy class of the matrix $g_{\sigma}$.
We have the following equality
$\langle 1_{\mathcal{O}_{1}}\ast\cdots\ast 1_{\mathcal{O}_{k}},1_{\{1\}}\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}=\dfrac{(q-1)}{2^{m}}\sum_{z\in\mathbb{F}_{q}^{*}}\langle 1_{O_{1}}\ast\cdots\ast 1_{O_{k}},1_{\{z\}}\rangle_{\operatorname{GL}_{2}(\mathbb{F}_{q})}.$ (5.6.5)
Indeed, the LHS of eq.(5.6.5) is
|$\dfrac{\Big{|}\{(x_{1},\dots,x_{k})\in\mathcal{O}_{1}\times\cdots\times\mathcal{O}_{k}\ |\ x_{1}\cdots x_{k}=1\}\Big{|}}{|\operatorname{PGL}_{2}(\mathbb{F}_{q})|}$
and the RHS of eq.(5.6.5) is
$\dfrac{\Bigg{|}\displaystyle\bigsqcup_{z\in\mathbb{F}_{q}^{*}}\{(y_{1},\dots,y_{k})\in O_{1}\times\cdots\times O_{k}\ |\ y_{1}\cdots y_{k}=z\}\Bigg{|}}{2^{m}|\operatorname{PGL}_{2}(\mathbb{F}_{q})|}.$
A similar argument to the one used in Proposition 3.2.1 in the geometric context, shows now that the projection map
$\bigsqcup_{z\in\mathbb{F}_{q}^{*}}\{(y_{1},\dots,y_{k})\in O_{1}\times\cdots\times O_{k}\ |\ y_{1}\cdots y_{k}=z\}\to\{(x_{1},\dots,x_{k})\in\mathcal{O}_{1}\times\cdots\times\mathcal{O}_{k}\ |\ x_{1}\cdots x_{k}=1\}$
$(y_{1},\dots,y_{k})\to(p(y_{1}),\dots,p(y_{k}))$
is a quotient map for the free $(\mathbb{Z}/2\mathbb{Z})^{m}$-action obtained by multiplication on the first $m$ coordinates.
Notice that $\langle 1_{O_{1}}\ast\cdots\ast 1_{O_{k}},1_{\{z\}}\rangle_{\operatorname{GL}_{2}(\mathbb{F}_{q})}=0$ if $\det(O_{1})\cdots\det(O_{k})\neq z^{2}$. In particular, we deduce that
$\det(O_{1})\cdots\det(O_{k})\not\in(\mathbb{F}_{q}^{*})^{2}\Rightarrow\langle 1_{\mathcal{O}_{1}}\ast\cdots\ast 1_{\mathcal{O}_{k}},1_{\{1\}}\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}=0$ (5.6.6)
where $(\mathbb{F}_{q}^{*})^{2}$ is the subgroup of squares and otherwise, if
$\det(O_{1})\cdots\det(O_{k})=\lambda_{O}^{2}$
with $\lambda_{O}\in\mathbb{F}_{q}^{*}$, then
$\langle 1_{\mathcal{O}_{1}}\ast\cdots\ast 1_{\mathcal{O}_{k}},1_{\{1\}}\rangle_{\operatorname{PGL}_{2}(\mathbb{F}_{q})}=\dfrac{(q-1)\left(\langle 1_{O_{1}}\ast\cdots\ast 1_{O_{k}},1_{\{\lambda_{O}\}}\rangle_{\operatorname{GL}_{2}(\mathbb{F}_{q})}+\langle 1_{O_{1}}\ast\cdots\ast 1_{O_{k}},1_{\{-\lambda_{O}\}}\rangle_{\operatorname{GL}_{2}(\mathbb{F}_{q})}\right)}{2^{m-1}}$ (5.6.7)
We apply Formula (5.6.7) in the following way. Recall that by hypothesis we are assuming that
$\det(C_{1})\cdots\det(C_{k})=(-1)^{m}\det(C_{m+1})\cdots\det(C_{k})\in(\mathbb{F}_{q}^{*})^{2}.$
Fix a subset $B\subseteq\{2,\dots,m\}$ and the corresponding class functions $\bf{Y}_{1},\dots,{\bf{Y}}_{k}$ as in section 5.1. Put $m_{1}=|B|+1$ and $m_{2}=m-m_{1}$. Since the convolution product is commutative, we have
$\left\langle{\bf Y}_{1}\ast\cdots\ast{\bf Y}_{k},1_{\{1\}}\right\rangle=\left\langle Y_{\epsilon}^{\ast m_{1}}\ast Y_{1}^{\ast m_{2}}\ast 1_{\mathcal{C}_{m+1}^{F}}\ast\cdots\ast 1_{\mathcal{C}_{k}^{F}},1_{\{1\}}\right\rangle=$ (5.6.8)
$\left\langle(1_{\mathcal{O}_{e}}-1_{\mathcal{O}_{\sigma}})^{\ast m_{1}}\ast(1_{\mathcal{O}_{e}}+1_{\mathcal{O}_{\sigma}})^{\ast m_{2}}\ast 1_{\mathcal{C}_{m+1}^{F}}\ast\cdots\ast 1_{\mathcal{C}_{k}^{F}},1_{\{1\}}\right\rangle=$ (5.6.9)
$\sum_{r=0}^{m}C_{|B|+1,m-|B|-1,r}\left\langle 1_{\mathcal{O}_{\sigma}}^{\ast r}\ast 1_{\mathcal{O}_{e}}^{\ast m-r}\ast 1_{\mathcal{C}_{m+1}^{F}}\ast\cdots\ast 1_{\mathcal{C}_{k}^{F}},1_{\{1\}}\right\rangle.$ (5.6.10)
For each $r$, we have that
$\det(O_{\sigma})^{r}\det(O_{e})^{m-r}\det(C_{m+1})\cdots\det(C_{k})=x^{2r}(-1)^{m}\det(C_{m+1})\cdots\det(C_{k}).$
Since $x^{2}\in\mathbb{F}_{q}^{*}\setminus(\mathbb{F}_{q}^{*})^{2}$, from (5.6.6), we have
$\left\langle 1_{\mathcal{O}_{\sigma}}^{\ast r}\ast 1_{\mathcal{O}_{e}}^{\ast m-r}\ast 1_{\mathcal{C}_{m+1}^{F}}\ast\cdots\ast 1_{\mathcal{C}_{k}^{F}},1_{\{1\}}\right\rangle=0,$
unless $r$ is even. Given the genericity of $\mathcal{C}$, Formula (5.6.7) and an argument similar to the one used at the end of the proof of Theorem 2.3.6 show that, for each $r$ even, we have
$\left\langle 1_{\mathcal{O}_{\sigma}}^{\ast r}\ast 1_{\mathcal{O}_{e}}^{\ast m-r}\ast 1_{\mathcal{C}_{m+1}^{F}}\ast\cdots\ast 1_{\mathcal{C}_{k}^{F}},1_{\{1\}}\right\rangle=\dfrac{1}{2^{m-1}}\mathbb{A}_{r}\left(\frac{1}{\sqrt{q}},\sqrt{q}\right).$
Theorem 5.1.3 is now a consequence of Theorem 3.4.5.